GNU MDK

GNU MIX Development Kit

Edition 1.2.9, for GNU mdk Version 1.2.9 November, 2015

by Jose Antonio Ortega Ruiz (jao@gnu.org)

This manual is for GNU MDK (version 1.2.9, November, 2015), a set of utilities for developing programs using Donald Knuth’s MIX mythical computer and MIXAL, its assembly language. Copyright \begin{matrix}c \\ ◯ \end{matrix} 2000, 2001, 2002, 2003, 2004, 2005, 2006, 2008, 2009, 2010, 2013, 2014, 2015 Free Software Foundation, Inc.

Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.2 or any later version published by the Free Software Foundation; with the Invariant Sections being “GNU General Public License”, with the Front-Cover Texts being “A GNU Manual,” and with the Back-Cover Texts as in (a) below. A copy of the license is included in the section entitled “GNU Free Documentation License”.

(a) The FSF’s Back-Cover Text is: “You have the freedom to copy and modify this GNU manual. Buying copies from the FSF supports it in developing GNU and promoting software freedom.”

Short Contents

Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Acknowledgements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1 Installing mdk. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2 MIX and MIXAL tutorial . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 3 Getting started . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 4 Emacs tools. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 5 mixasm, the MIXAL assembler. . . . . . . . . . . . . . . . . . . . . . . . . . 45 6 mixvm, the MIX computer simulator . . . . . . . . . . . . . . . . . . . . . 47 7 gmixvm, the GTK virtual machine . . . . . . . . . . . . . . . . . . . . . . . 59 8 mixguile, the Scheme virtual machine . . . . . . . . . . . . . . . . . . . 65 9 Reporting Bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 A Copying . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 Instructions and commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95

Table of Contents

Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

Acknowledgements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

1 Installing mdk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

1.1 Download the source tarball . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 1.2 Requirements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 1.3 Basic installation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 1.4 Emacs support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 1.5 Special configure flags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 1.6 Supported platforms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

2 MIX and MIXAL tutorial. . . . . . . . . . . . . . . . . . . . . . 9

2.1 The MIX computer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.1.1 MIX architecture . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.1.2 MIX instruction set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

2.1.2.1 Instruction structure . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 2.1.2.2 Loading operators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 2.1.2.3 Storing operators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 2.1.2.4 Arithmetic operators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 2.1.2.5 Address transfer operators . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 2.1.2.6 Comparison operators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 2.1.2.7 Jump operators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 2.1.2.8 Input-output operators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 2.1.2.9 Conversion operators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 2.1.2.10 Shift operators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 2.1.2.11 Miscellaneous operators . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 2.1.2.12 Execution times . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

2.2 MIXAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 2.2.1 Basic program structure . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 2.2.2 MIXAL directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 2.2.3 Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 2.2.4 W-expressions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 2.2.5 Local symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 2.2.6 Literal constants . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

3 Getting started . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27

3.1 Writing a source file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 3.2 Compiling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 3.3 Running the program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

3.3.1 Non-interactive mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 3.3.2 Interactive mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

3.3.3 Debugging commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 3.4 Using mixguile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

3.4.1 The mixguile shell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 3.4.2 Additional MIX Scheme functions . . . . . . . . . . . . . . . . . . . . . . . . 34 3.4.3 Defining new functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 3.4.4 Hook functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36

3.4.4.1 Command hooks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 3.4.4.2 Break hooks. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39

3.4.5 Scheme scripts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 3.5 Using Scheme in mixvm and gmixvm . . . . . . . . . . . . . . . . . . . . . . . . . . . 40

4 Emacs tools . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

4.1 MIXAL mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 4.1.1 Basics. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 4.1.2 Help system. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 4.1.3 Compiling and running . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44

4.2 GUD integration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44

5 mixasm, the MIXAL assembler . . . . . . . . . . . . . . . . 45

5.1 Invoking mixasm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

6 mixvm, the MIX computer simulator . . . . . . . . . 47

6.1 Invoking mixvm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 6.2 Interactive commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48

6.2.1 File commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 6.2.2 Debug commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 6.2.3 State commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 6.2.4 Configuration commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 6.2.5 Scheme commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57

6.3 MIX block devices . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57

7 gmixvm, the GTK virtual machine . . . . . . . . . . . . 59

7.1 Invoking gmixvm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 7.2 MIXVM console . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 7.3 MIX virtual machine . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 7.4 MIXAL source view . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 7.5 MIX devices view . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 7.6 Menu and status bars . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61

8 mixguile, the Scheme virtual machine . . . . . . . 65

8.1 Invoking mixguile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 8.2 Scheme functions reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66

8.2.1 mixvm command wrappers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 8.2.2 Hook functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 8.2.3 Additional VM functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69

9 Reporting Bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73

Appendix A Copying . . . . . . . . . . . . . . . . . . . . . . . . . . . 75

A.1 GNU General Public License . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 A.2 GNU Free Documentation License . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86

Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93

Instructions and commands . . . . . . . . . . . . . . . . . . . . . . 95

Introduction

In his book series The Art of Computer Programming (published by Addison Wesley), D. Knuth uses an imaginary computer, the MIX, and its associated machine-code and assembly languages to illustrate the concepts and algorithms as they are presented.

The MIX’s architecture is a simplified version of those found in real CISC CPUs, and the MIX assembly language (MIXAL) provides a set of primitives that will be very familiar to any person with a minimum experience in assembly programming. The MIX/MIXAL definition is powerful and complete enough to provide a virtual development platform for writing quite complex programs, and close enough to real computers to be worth using when learning programming techniques. At any rate, if you want to learn or improve your programming skills, a MIX development environment would come in handy.

The mdk package aims at providing such virtual development environment on a GNU box. Thus, mdk offers you a set of utilities to simulate the MIX computer and to write, compile, run and debug MIXAL programs. As of version 1.2.9, mdk includes the following programs:

mixasm MIXAL assembler. Assembler which translates MIXAL source files into programs that can be run (and debugged) by mixvm, mixguile or gmixvm.

mixvm MIX virtual machine. Emulation of the MIX computer with a CLI.

gmixvm A GTK+ GUI for the MIX virtual machine. Provides all of mixvm functionality accessible through a graphical interface.

mixguile A Guile shell, with an embedded MIX virtual machine and built-in commands to manipulate it using Scheme.

mixal-mode.el

An Emacs major mode for MIXAL source files editing, providing syntax highlighting, documentation lookup and invocation of mixvm within Emacs.

mixvm.el This elisp program allows running mixvm inside an Emacs GUD buffer, providing concurrent edition and debugging of MIXAL programs.

mixvm and gmixvm implement a simulator of the MIX computer, giving you a virtual machine for executing and debugging MIX programs. These binary programs could be written by hand, but it is easier to produce them compiling MIXAL source files, using the MIXAL assembler mixasm. On the other hand, mixguile offers you the possibility of manipulating a MIX virtual machine through a set of Scheme functions, so that you can use this programming language to interact with the virtual machine. In addition, mixvm and gmixvm are also able to interpret Scheme scripts (using an embedded Guile interpreter), that is, you can use Scheme as an extension language to add new functionalities to these programs.

This manual gives you a tutorial of MIX and MIXAL, and a thorough description of the use of the mdk utilities.

Acknowledgements

Many people have further contributed to mdk by reporting problems, suggesting various improvements, or submitting actual code. Here is a list of these people. Please, help me keep it complete and exempt of errors.

• Philip Ellis King provided MIXAL test programs pinpointing bugs in the first mdk release, and useful discussions as well. Philip has also contributed with the Emacs port of mixvm and influenced the gmixvm GUI design with insightful comments and prototypes.

• Aleix Conchillo has been following mdk’s development for many years, indefatigably chasing and fixing bugs, and suggesting many improvements. He’s also the original author of the Fink and Macports ports.

• Pieter E J Pareit is the author of the Emacs MIXAL mode, and has also contributed many bug fixes.

• Michael Scholz is the author of the German translation of mdk’s user interface.

• Sergey Poznyakoff provided patches to mixlib/mix scanner.l improving MIXAL compliance.

• Sergey Litvin implemented the instructions SLB, SRB, JAE, JAO, JXE, and JXO from volume 2 of TAOCP.

• Francesc Xavier Noria kindly and thoroughly reviewed the mdk documentation, providing insightful advice.

• Eric S. Raymond contributed the documentation file MIX.DOC and the samples

elevator.mixal and mistery.mixal from his MIXAL package.

• Nelson H. F. Beebe has tested mdk in a lot of Unix platforms, suggesting portability enhancements to the source code.

• Ryan Schmidt, Agustin Navarro, Ying-Chieh Liao, Adrian Bunk, Baruch Even, and Ronald Cole ported mdk to different platforms, and created and/or maintain packages for it.

• Jason Uhlenkott, Andrew Hood, Radu Butnaru, Ruslan Batdalov, WeiZheng, Sascha Wilde, Michael Vernov and Xiaofeng Zhao reported bugs and suggested fixes to them.

• Joshua Davies, Eli Bendersky, Milan Bella and Jens Seidel reported bugs on the documentation.

• Christoph von Nathusius, Stephen Ramsay and Johan Swanljung tested mdk on different platforms, and helped fixing the configuration process in them.

• Richard Stallman suggested various improvements to the documentation and has always kept an eye on each mdk release.

• mdk was inspired by Darius Bacon’s MIXAL program (http://www.accesscom.com/

~darius/).

1 Installing mdk

1.1 Download the source tarball

GNU mdk is distributed as a source tarball available for download in the following URLs:

  • ftp://ftp.gnu.org/pub/gnu/mdk
  • GNU mirrors (http://www.gnu.org/prep/ftp.html)

The above sites contain the latest stable releases of mdk. The development branch is available as a Git (http://git-scm.com/) repository located at^{1}

  • git://git.savannah.gnu.org/mdk.git

After you have downloaded the source tarball, unpack it in a directory of your choice using the command:

     tar xfvz mdk-X.Y.tar.gz

where X.Y stands for the downloaded version (the current stable release being version 1.2.9).

1.2 Requirements

In order to build and install mdk, you will need the following libraries installed in your system:

− GLIB 2.16.0 (http://www.gtk.org) (required)

 − GNU Flex 2.5 (http://www.gnu.org/software/flex/flex.html) (required)

− GTK 2.16.0 (http://www.gtk.org) (optional)

 − Libglade 2.6.0 (http://ftp.gnome.org/pub/GNOME/sources/libglade/2.6/) (op-

tional)

 − GNU Readline (http://cnswww.cns.cwru.edu/php/chet/readline/rltop.html)

(optional)

 − GNU Libguile 2.0 (http://www.gnu.org/software/guile) (optional)

If present, readline and history are used to provide command completion and history management to the command line MIX virtual machine, mixvm. GTK+ and libglade are needed if you want to build the graphical interface to the MIX virtual machine, gmixvm. Finally, if libguile is found, the mdk utilities will be compiled with Guile support and will be extensible using Scheme.

Please note: you need both the libraries and the headers; this means both the library package and the -dev package if you do not compile your libraries yourself (ex: installing

libgtk2.0-0 and libgtk2.0-0-dev on Debian).

1 See mdk’s Git page (https://savannah.gnu.org/git/?group=mdk) for more information on using the

unstable source tree. Note, however, that the rest of this manual is about the stable release.

1.3 Basic installation

mdk uses GNU Autoconf and Automake tools, and, therefore, should be built and installed without hassle using the following commands inside the source directory:

     ./configure
     make
     make install

where the last one must be run as root.

The first command, configure, will setup the makefiles for your system. In particular, configure will look for GTK+ and libglade, and, if they are present, will generate the appropriate makefiles for building the gmixvm graphical user interface. Upon completion, you should see a message with the configuration results like the following:

     *** GNU MDK 1.2.9 has been successfully configured. ***
     Type 'make' to build the following utilities:
         - mixasm (MIX assembler)
         - mixvm (MIX virtual machine, with readline support,
                  with guile support)
         - gmixvm (mixvm GTK+ GUI, with guile support)
         - mixguile (the mixvm guile shell)

where the last lines may be missing if you lack the above mentioned libraries.

The next command, make, will actually build the mdk programs in the following locations:

 − mixutils/mixasm
 − mixutils/mixvm
 − mixgtk/gmixvm
 − mixguile/mixguile

You can run these programs from within their directories, but I recommend you to install them in proper locations using make install from a root shell.

1.4 Emacs support

mdk includes extensive support for Emacs. Upon installation, all the elisp code is installed in PREFIX/share/mdk, where PREFIX stands for your installation root directory (e.g. /usr/local). You can copy the elisp files to a directory that is in your load-path, or you can add the above directory to it. Assuming that the installing prefix is /usr/local, you can do it by adding to your .emacs file the following line:

     (setq load-path (cons "/usr/local/share/mdk" load-path))

MIXAL programs can be written using Emacs and the elisp program share/mdk/mixalmode.el, contributed by Pieter E. J. Pareit. It provides font locking, interactive help, compiling assistance and invocation of the MIX virtual machine via a new major mode called mixal-mode. To start mixal-mode automatically whenever you edit a MIXAL source file, add the following lines to your .emacs file:

     (autoload 'mixal-mode "mixal-mode" t)

(add-to-list 'auto-mode-alist '("\\.mixal\\'" . mixal-mode))

In addition, mixvm can be run within an Emacs GUD buffer using the elisp program share/mdk/mixvm.el, contributed by Philip E. King. mixvm.el provides an interface between mdk’s mixvm and Emacs, via GUD. Place this file in your load-path, optionally adding the following line to your .emacs file:

(autoload 'mixvm "mixvm" "mixvm/gud interaction" t)

1.5 Special configure flags

You can fine-tune the configuration process using the following switches with configure:

--enable-gui[=yes|no]  [User Option]
[User Option] --disable-gui

Enables/disables the build of the MIX virtual machine GUI (gmixvm). If the required libraries are missing (see Section 1.2 [Requirements], page 5) the configure script with automatically disable this feature.

--with-guile[=yes|no]  [User Option]
[User Option] --without-guile

Enables/disables the Guile support for mixvm and gmixvm, and the build of mixguile. If the required libraries are missing (see Section 1.2 [Requirements], page 5) the configure script with automatically disable this feature.

--with-readline[=yes|no]  [User Option]
[User Option] --without-readline

Enables/disables the GNU Readline support for mixvm. If the required libraries are missing (see Section 1.2 [Requirements], page 5) the configure script with automatically disable this feature.

For additional, boilerplate configure options, see the INSTALL file, or run

configure --help

1.6 Supported platforms

GNU MDK has been tested in the following platforms:

• Debian GNU/Linux 2.2, 2.3, 3.0, 3.1, 3.2, 4.0, 5.0, 6.0, sid

• Redhat GNU/Linux 8.0 (Ronald Cole), 7.0 (Agustin Navarro), 6.2 (Roberto Ferrero)

• Mandrake 8.0 (Agustin Navarro)

• FreeBSD 4.2, 4.3, 4.4, 4.5 (Ying-Chieh Liao), 5.2

• Solaris 2.8/gcc 2.95.3 (Stephen Ramsay)

• MS Windows 98 SE/Cygwin 1.1.8-2 (Christoph von Nathusius)^{2}

• Mac OS X 10.1.2 (Johan Swanljung), Mac OS X 10.4.x, 10.5 (Darwin Port by Aleix Conchillo).

2 Caveats: Christoph has only tested mixvm and mixasm on this platform, using gcc 2.95.3-2, GLIB 1.2.10

and GNU readline 4.1-2. He has reported missing history functionalities on a first try. If you find problems with history/readline functionality, please try a newer/manually installed readline version.

• AMD Athlon, GNU/Linux version 2.4.2-2smp (Red Hat 7.1 (Seawolf)) (N. H. F. Beebe)

• Apple PowerPC G3, GNU/Linux 2.2.18-4hpmac (Red Hat Linux/PPC 2000 Q4) (N. H. F. Beebe)

• DEC Alpha, GNU/Linux 2.2.19-6.2.1 (Red Hat 6.2) (N. H. F. Beebe)

• Compaq/DEC Alpha OSF/1 4.0F [ONLY after adding rsync’s snprintf() implementation] (N. H. F. Beebe)

• IBM PowerPC AIX 4.2 (N. H. F. Beebe)

• Intel Pentium III, GNU/Linux 2.4.9-31smp (Red Hat 7.2 (Enigma)) (N. H. F. Beebe)

• SGI Origin 200, IRIX 6.5 (N. H. F. Beebe)

• Sun SPARC, GNU/Linux 2.2.19-6.2.1 (Red Hat 6.2) (N. H. F. Beebe)

• Sun SPARC, Solaris 2.8 (N. H. F. Beebe) mdk will probably work on any GNU/Linux or BSD platform. If you try it in a platform

not listed above, please send a mail to the author (jao@gnu.org).

2 MIX and MIXAL tutorial

In the book series The Art of Computer Programming, by D. Knuth, a virtual computer, the MIX, is used by the author (together with the set of binary instructions that the virtual CPU accepts) to illustrate the algorithms and skills that every serious programmer should master. Like any other real computer, there is a symbolic assembler language that can be used to program the MIX: the MIX assembly language, or MIXAL for short. In the following subsections you will find a tutorial on these topics, which will teach you the basics of the MIX architecture and how to program a MIX computer using MIXAL.

2.1 The MIX computer

In this section, you will find a description of the MIX computer, its components and instruction set.

2.1.1 MIX architecture

The basic information storage unit in the MIX computer is the byte, which stores positive values in the range 0-63 . Note that a MIX byte can be then represented as 6 bits, instead of the common 8 bits for a regular byte. Unless otherwise stated, we shall use the word byte to refer to a MIX 6-bit byte.

A MIX word is defined as a set of 5 bytes plus a sign. The bytes within a word are numbered from 1 to 5, being byte number one the most significant one. The sign is denoted by index 0. Graphically,

      -----------------------------------------------
     |   0   |   1   |   2   |   3   |   4   |   5   |
      -----------------------------------------------
     |  +/-  | byte  | byte  | byte  | byte  | byte  |
      -----------------------------------------------

Sample MIX words are ‘- 12 00 11 01 63’ and ‘+ 12 11 34 43 00’.

You can refer to subfields within a word using a field specification or fspec of the form “(L:R)”, where L denotes the first byte, and R the last byte of the subfield. When L is zero, the subfield includes the word’s sign. An fspec can also be represented as a single value F, given by F = 8*L + R (thus the fspec ‘(1:3)’, denoting the first three bytes of a word, is represented by the integer 11).

The MIX computer stores information in registers, that can store either a word or two bytes and sign (see below), and memory cells, each one containing a word. Specifically, the MIX computer has 4000 memory cells with addresses 0 to 3999 (i.e., two bytes are enough to address a memory cell) and the following registers:

rA A register. General purpose register holding a word. Usually its contents serves as the operand of arithmetic and storing instructions.

rX X register. General purpose register holding a word. Often it acts as an extension or a replacement of ‘rA’.

rJ J (jump) register. This register stores positive two-byte values, usually representing a jump address.

rI1, rI2, rI3, rI4, rI5, rI6

Index registers. These six registers can store a signed two-byte value. Their contents are used as indexing values for the computation of effective memory addresses.

In addition, the MIX computer contains: − An overflow toggle (a single bit with values on or off ). In this manual, this toggle is

denoted ov. − A comparison indicator (having three values: EQUAL, GREATER or LESS). In this

manual, this indicator is denoted cm, and its possible values are abbreviated as E, G and L.

− Input-output block devices. Each device is labelled as un, where n runs from 0 to 20. In Knuth’s definition, u0 through u7 are magnetic tape units, u8 through 15 are disks and drums, u16 is a card reader, u17 is a card writer, u18 is a line printer and, u19 is a typewriter terminal, and u20, a paper tape. Our implementation maps these devices to disk files, except for u19, which represents the standard output.

As noted above, the MIX computer communicates with the external world by a set of input-output devices which can be “connected” to it. The computer interchanges information using blocks of words whose length depends on the device at hand (see Section 6.3 [Devices], page 57). These words are interpreted by the device either as binary information (for devices 0-16), or as representing printable characters (devices 17-20). In the last case, each MIX byte is mapped onto a character according to the following table:

00  01 A 02 B 03 C
04 D 05 E 06 F 07 G
08 H 09 I 10 ~ 11 J
12 K 13 L 14 M 15 N
16 O 17 P 18 Q 19 R
20 [ 21 # 22 S 23 T
24 U 25 V 26 W 27 X
28 Y 29 Z 30 0 31 1
32 2 33 3 34 4 35 5
36 6 37 7 38 8 39 9
40 . 41 , 42 ( 43 )
44 + 45 - 46 * 47 /
48 = 49 $ 50 < 51 >
52 @ 53 ; 54 : 55 ’

The value 0 represents a whitespace. The characters ~, [ and # correspond to symbols not representable as ASCII characters (uppercase delta, sigma and gamma, respectively), and byte values 56-63 have no associated character.

Finally, the MIX computer features a virtual CPU which controls the above components, and which is able to execute a rich set of instructions (constituting its machine language, similar to those commonly found in real CPUs), including arithmetic, logical, storing, comparison and jump instructions. Being a typical von Neumann computer, the MIX CPU fetches binary instructions from memory sequentially (unless a jump instruction is found), and stores the address of the next instruction to be executed in an internal register called location counter (also known as program counter in other architectures).

The next section, See Section 2.1.2 [MIX instruction set], page 11, gives a complete description of the available MIX binary instructions.

2.1.2 MIX instruction set

The following subsections fully describe the instruction set of the MIX computer. We begin with a description of the structure of binary instructions and the notation used to refer to their subfields. The remaining subsections are devoted to describing the actual instructions available to the MIX programmer.

2.1.2.1 Instruction structure

MIX instructions are codified as words with the following subfield structure:

Subfield fspec Description

ADDRESS (0:2)

The first two bytes plus sign are the address field. Combined with the INDEX field, denotes the memory address to be used by the instruction.

INDEX (3:3)

The third byte is the index, normally used for indexing the address^{1}.

MOD (4:4)

Byte four is used either as an operation code modifier or as a field specification.

OPCODE (5:5)

The last (least significant) byte in the word denotes the operation code.

or, graphically,

      ------------------------------------------------
     |   0   |   1   |   2   |   3   |   4   |   5    |
      ------------------------------------------------
     |        ADDRESS        | INDEX |  MOD  | OPCODE |
      ------------------------------------------------

For a given instruction, ‘M’ stands for the memory address obtained after indexing the ADDRESS subfield (using its INDEX byte), and ‘V’ is the contents of the subfield indicated by MOD of the memory cell with address ‘M’. For instance, suppose that we have the following contents of MIX registers and memory cells:

     [rI2] = + 00 63
     [31] = - 10 11 00 11 22

where ‘[n]’ denotes the contents of the nth memory cell and ‘[rI2]’ the contents of register ‘rI2’^{2}. Let us consider the binary instruction ‘I = - 00 32 02 11 10’. For this instruction we have:

     ADDRESS = - 00 32 = -32
     INDEX = 02 = 2
     MOD = 11 = (1:3)
     OPCODE = 10

1 The actual memory address the instruction refers to, is obtained by adding to ADDRESS the value of

the ‘rI’ register denoted by INDEX.

2 In general, ‘[X]’ will denote the contents of entity ‘X’; thus, by definition, ‘V = [M](MOD)’.

     M = ADDRESS + [rI2] = -32 + 63 = 31
     V = [M](MOD) = (- 10 11 00 11 22)(1:3) = + 00 00 10 11 00

Note that, when computing ‘V’ using a word and an fspec, we apply a left padding to the bytes selected by ‘MOD’ to obtain a complete word as the result.

In the following subsections, we will assign to each MIX instruction a mnemonic, or symbolic name. For instance, the mnemonic of ‘OPCODE’ 10 is ‘LD2’. Thus we can rewrite the above instruction as

     LD2  -32,2(1:3)

or, for a generic instruction:

     MNEMONIC  ADDRESS,INDEX(MOD)

Some instructions are identified by both the OPCODE and the MOD fields. In these cases, the MOD will not appear in the above symbolic representation. Also when ADDRESS or INDEX are zero, they can be omitted. Finally, MOD defaults to (0:5) (meaning the whole word).

2.1.2.2 Loading operators

The following instructions are used to load memory contents into a register.

LDA Put in rA the contents of cell no. M. OPCODE = 8, MOD = fspec. rA <- V.

LDX Put in rX the contents of cell no. M. OPCODE = 15, MOD = fspec. rX <- V.

LDi Put in rIi the contents of cell no. M. OPCODE = 8 + i, MOD = fspec. rIi <-

          V.

LDAN Put in rA the contents of cell no. M, with opposite sign. OPCODE = 16, MOD = fspec. rA <- -V.

LDXN Put in rX the contents of cell no. M, with opposite sign. OPCODE = 23, MOD = fspec. rX <- -V.

LDiN Put in rIi the contents of cell no. M, with opposite sign. OPCODE = 16 + i, MOD = fspec. rIi <- -V.

In all the above load instructions the ‘MOD’ field selects the bytes of the memory cell with address ‘M’ which are loaded into the requisite register (indicated by the ‘OPCODE’). For instance, the word ‘+ 00 13 01 27 11’ represents the instruction

     LD3    13,1(3:3)
      ^^^^
      |      | |  |
      |      | |   --- MOD = 27 = 3*8 + 3
      |      |  --- INDEX = 1
      |       --- ADDRESS = 00 13
       --- OPCODE = 11

Let us suppose that, prior to this instruction execution, the state of the MIX computer is the following:

     [rI1] = - 00 01
     [rI3] = + 24 12
     [12] = - 01 02 03 04 05

As, in this case, ‘M = 13 + [rI1] = 12’, we have

     V = [M](3:3) = (- 01 02 03 04 05)(3:3)
       = + 00 00 00 00 03

(note that the specified subfield is left-padded with null bytes to complete a word). Hence, the MIX state, after the instruction execution, will be

     [rI1] = - 00 01
     [rI3] = + 00 03
     [12] = - 01 02 03 04 05

To further illustrate loading operators, the following table shows the contents of ‘rX’ after different ‘LDX’ instructions:

‘LDX 12(0:0) [rX] = - 00 00 00 00 00’
‘LDX 12(0:1) [rX] = - 00 00 00 00 01’
‘LDX 12(3:5) [rX] = + 00 00 03 04 05’
‘LDX 12(3:4) [rX] = + 00 00 00 03 04’
‘LDX 12(0:5) [rX] = - 01 02 03 04 05’

2.1.2.3 Storing operators

The following instructions are the inverse of the load operations: they are used to store a subfield of a register into a memory location. Here, MOD represents the subfield of the memory cell that is to be overwritten with bytes from a register. These bytes are taken beginning by the rightmost side of the register.

STA Store rA. OPCODE = 24, MOD = fspec. V <- rA.

STX Store rX. OPCODE = 31, MOD = fspec. V <- rX.

STi Store rIi. OPCODE = 24 + i, MOD = fspec. V <- rIi.

STJ Store rJ. OPCODE = 32, MOD = fspec. V <- rJ.

STZ Store zero. OPCODE = 33, MOD = fspec. V <- 0.

By way of example, consider the instruction ‘STA 1200(2:3)’. It causes the MIX to fetch bytes no. 4 and 5 of register A and copy them to bytes 2 and 3 of memory cell no. 1200 (remember that, for these instructions, MOD specifies a subfield of the memory address). The other bytes of the memory cell retain their values. Thus, if prior to the instruction execution we have

     [1200] = - 20 21 22 23 24
     [rA] = + 01 02 03 04 05

we will end up with

     [1200] = - 20 04 05 23 24
     [rA] = + 01 02 03 04 05

As a second example, ‘ST2 1000(0)’ will set the sign of ‘[1000]’ to that of ‘[rI2]’.

2.1.2.4 Arithmetic operators

The following instructions perform arithmetic operations between rA and rX register and memory contents.

ADD Add and set OV if overflow. OPCODE = 1, MOD = fspec. rA <- rA +V.

SUB Sub and set OV if overflow. OPCODE = 2, MOD = fspec. rA <- rA - V.

MUL Multiply V times rA and store the 10-bytes product in rAX. OPCODE = 3, MOD = fspec. rAX <- rA x V.

DIV rAX is considered a 10-bytes number, and it is divided by V. OPCODE = 4, MOD = fspec. rA <- rAX / V, rX <- reminder.

In all the above instructions, ‘[rA]’ is one of the operands of the binary arithmetic operation, the other being ‘V’ (that is, the specified subfield of the memory cell with address ‘M’), padded with zero bytes on its left-side to complete a word. In multiplication and division, the register ‘X’ comes into play as a right-extension of the register ‘A’, so that we are able to handle 10-byte numbers whose more significant bytes are those of ‘rA’ (the sign of this 10-byte number is that of ‘rA’: ‘rX’’s sign is ignored).

Addition and subtraction of MIX words can give rise to overflows, since the result is stored in a register with room to only 5 bytes (plus sign). When this occurs, the operation result modulo 1,073,741,823 (the maximum value storable in a MIX word) is stored in ‘rA’, and the overflow toggle is set to TRUE.

2.1.2.5 Address transfer operators

In these instructions, ‘M’ (the address of the instruction after indexing) is used as a number instead of as the address of a memory cell. Consequently, ‘M’ can have any valid word value (i.e., it’s not limited to the 0-3999 range of a memory address).

ENTA Enter ‘M’ in [rA]. OPCODE = 48, MOD = 2. rA <- M.

ENTX Enter ‘M’ in [rX]. OPCODE = 55, MOD = 2. rX <- M.

ENTi Enter ‘M’ in [rIi]. OPCODE = 48 + i, MOD = 2. rIi <- M.

ENNA Enter ‘-M’ in [rA]. OPCODE = 48, MOD = 3. rA <- -M.

ENNX Enter ‘-M’ in [rX]. OPCODE = 55, MOD = 3. rX <- -M.

ENNi Enter ‘-M’ in [rIi]. OPCODE = 48 + i, MOD = 3. rIi <- -M.

INCA Increase [rA] by ‘M’. OPCODE = 48, MOD = 0. rA <- rA + M.

INCX Increase [rX] by ‘M’. OPCODE = 55, MOD = 0. rX <- rX + M.

INCi Increase [rIi] by ‘M’. OPCODE = 48 + i, MOD = 0. rIi <- rIi + M.

DECA Decrease [rA] by ‘M’. OPCODE = 48, MOD = 1. rA <- rA - M.

DECX Decrease [rX] by ‘M’. OPCODE = 55, MOD = 1. rX <- rX - M.

DECi Decrease [rIi] by ‘M’. OPCODE = 48 + i, MaOD = 0. rIi <- rIi - M.

In the above instructions, the subfield ‘ADDRESS’ acts as an immediate (indexed) operand, and allow us to set directly the contents of the MIX registers without an indirection to the memory cells (in a real CPU this would mean that they are faster that the previously discussed instructions, whose operands are fetched from memory). So, if you want to store in ‘rA’ the value -2000 (- 00 00 00 31 16), you can use the binary instruction + 31 16 00 03 48, or, symbolically,

     ENNA 2000

Used in conjunction with the store operations (‘STA’, ‘STX’, etc.), these instructions also allow you to set memory cells contents to concrete values.

Note that in these address transfer operators, the ‘MOD’ field is not a subfield specificator, but serves to define (together with ‘OPCODE’) the concrete operation to be performed.

2.1.2.6 Comparison operators

So far, we have learned how to move values around between the MIX registers and its memory cells, and also how to perform arithmetic operations using these values. But, in order to write non-trivial programs, other functionalities are needed. One of the most common is the ability to compare two values, which, combined with jumps, will allow the execution of conditional statements. The following instructions compare the value of a register with ‘V’, and set the cm indicator to the result of the comparison (i.e. to ‘E’, ‘G’ or ‘L’, equal, greater or lesser respectively).

CMPA Compare [rA] with V. OPCODE = 56, MOD = fspec.

CMPX Compare [rX] with V. OPCODE = 63, MOD = fspec.

CMPi Compare [rIi] with V. OPCODE = 56 + i, MOD = fspec.

As explained above, these instructions modify the value of the MIX comparison indicator; but maybe you are asking yourself how do you use this value: enter jump operators, in the next subsection.

2.1.2.7 Jump operators

The MIX computer has an internal register, called the location counter, which stores the address of the next instruction to be fetched and executed by the virtual CPU. You cannot directly modify the contents of this internal register with a load instruction: after fetching the current instruction from memory, it is automatically increased in one unit by the MIX. However, there is a set of instructions (which we call jump instructions) which can alter the contents of the location counter provided some condition is met. When this occurs, the value of the next instruction address that would have been fetched in the absence of the jump is stored in ‘rJ’ (except for JSJ), and the location counter is set to the value of ‘M’ (so that the next instruction is fetched from this new address). Later on, you can return to the point when the jump occurred reading the address stored in ‘rJ’.

The MIX computer provides the following jump instructions: With these instructions you force a jump to the specified address. Use ‘JSJ’ if you do not care about the return address.

JMP Unconditional jump. OPCODE = 39, MOD = 0.

JSJ Unconditional jump, but rJ is not modified. OPCODE = 39, MOD = 1.

These instructions check the overflow toggle to decide whether to jump or not.

JOV Jump if OV is set (and turn it off). OPCODE = 39, MOD = 2.

JNOV Jump if OV is not set (and turn it off). OPCODE = 39, MOD = 3.

In the following instructions, the jump is conditioned to the contents of the comparison flag:

JL Jump if [CM] = L. OPCODE = 39, MOD = 4.

JE Jump if [CM] = E. OPCODE = 39, MOD = 5.

JG Jump if [CM] = G. OPCODE = 39, MOD = 6.

JGE Jump if [CM] does not equal L. OPCODE = 39, MOD = 7.

JNE Jump if [CM] does not equal E. OPCODE = 39, MOD = 8.

JLE Jump if [CM] does not equal G. OPCODE = 39, MOD = 9.

You can also jump conditioned to the value stored in the MIX registers, using the following instructions:

JAN
JAZ
JAP
JANN
JANZ
JANP
JAE

JAO Jump if the content of rA is, respectively, negative, zero, positive, non-negative, non-zero, non-positive, even or odd. OPCODE = 40, MOD = 0, 1, 2, 3, 4, 5, 6, 7.

JXN
JXZ
JXP
JXNN
JXNZ
JXNP
JXE

JXO Jump if the content of rX is, respectively, negative, zero, positive, non-negative, non-zero, non-positive, even or odd. OPCODE = 47, MOD = 0, 1, 2, 3, 4, 5, 6, 7.

JiN
JiZ
JiP
JiNN
JiNZ

JiNP Jump if the content of rIi is, respectively, negative, zero, positive, non-negative, non-zero or non-positive. OPCODE = 40 + i, MOD = 0, 1, 2, 3, 4, 5.

2.1.2.8 Input-output operators

As explained in previous sections (see Section 2.1.1 [MIX architecture], page 9), the MIX computer can interact with a series of block devices. To that end, you have at your disposal the following instructions:

IN Transfer a block of words from the specified unit to memory, starting at address M. OPCODE = 36, MOD = I/O unit.

OUT Transfer a block of words from memory (starting at address M) to the specified unit. OPCODE = 37, MOD = I/O unit.

IOC Perform a control operation (given by M) on the specified unit. OPCODE = 35, MOD = I/O unit.

JRED Jump to M if the specified unit is ready. OPCODE = 38, MOD = I/O unit.

JBUS Jump to M if the specified unit is busy. OPCODE = 34, MOD = I/O unit.

In all the above instructions, the ‘MOD’ subfile must be in the range 0-20, since it denotes the operation’s target device. The ‘IOC’ instruction only makes sense for tape devices (‘MOD’ = 0-7 or 20): it shifts the read/write pointer by the number of words given by ‘M’ (if it equals zero, the tape is rewound)^{3}.

2.1.2.9 Conversion operators

The following instructions convert between numerical values and their character representations.

NUM Convert rAX, assumed to contain a character representation of a number, to its numerical value and store it in rA. OPCODE = 5, MOD = 0.

CHAR Convert the number stored in rA to a character representation and store it in rAX. OPCODE = 5, MOD = 1.

Digits are represented in MIX by the range of values 30-39 (digits 0-9). Thus, if the contents of ‘rA’ and ‘rX’ are, for instance,

     [rA] = + 30 30 31 32 33
     [rX] = + 31 35 39 30 34

the represented number is 0012315904, and ‘NUM’ will store this value in ‘rA’ (i.e., we end up with ‘[rA]’ = + 0 46 62 52 0 = 12315904).

If any byte in ‘rA’ or ‘rB’ does not belong to the range 30-39, it is interpreted by ‘NUM’ as the digit obtained by taking its value modulo 10. E.g. values 0, 10, 20, 30, 40, 50, 60 all represent the digit 0; 2, 12, 22, etc. represent the digit 2, and so on. For instance, the number 0012315904 mentioned above could also be represented as

     [rA] = + 10 40 31 52 23
     [rX] = + 11 35 49 20 54

‘CHAR’ performs the inverse operation, using only the values 30 to 39 for representing digits 0-9.

2.1.2.10 Shift operators

The following instructions perform byte-wise shifts of the contents of ‘rA’ and ‘rX’.

3 In Knuth’s original definition, there are other control operations available, but they do not make sense

when implementing the block devices as disk files (as we do in mdk simulator). For the same reason,

mdk devices are always ready, since all input-output operations are performed using synchronous system

calls.

SLA
SRA
SLAX
SRAX
SLC

SRC Shift rA or rAX left, right, or rAX circularly (see example below) left or right. M specifies the number of bytes to be shifted. OPCODE = 6, MOD = 0, 1, 2, 3, 4, 5.

The following instructions perform binary shifts of the contents of ‘rA’ and ‘rX’.

SLB

SRB Shift rAX left or right binary. M specifies the number of binary places to shift. OPCODE = 6, MOD = 6, 7

If we begin with, say, ‘[rA]’ = - 01 02 03 04 05, we would have the following modifications to ‘rA’ contents when performing the instructions on the left column:

SLA 2 [rA] = - 03 04 05 00 00
SLA 6 [rA] = - 00 00 00 00 00
SRA 1 [rA] = - 00 01 02 03 04

Note that the sign is unaffected by shift operations. On the other hand, ‘SLC’, ‘SRC’, ‘SLAX’, ‘SRAX’, ‘SLB’ and ‘SRB’ treat ‘rA’ and ‘rX’ as a single 10-bytes register (ignoring again the signs). For instance, if we begin with ‘[rA]’ = + 01 02 03 04 05 and ‘[rX]’ = - 06 07 08 09 10, we would have:

SLC 3 [rA] = + 04 05 06 07 08 [rX] = - 09 10 01 02 03
SLAX 3 [rA] = + 04 05 06 07 08 [rX] = - 09 10 00 00 00
SRC 4 [rA] = + 07 08 09 10 01 [rX] = - 02 03 04 05 06
SRAX 4 [rA] = + 00 00 00 00 01 [rX] = - 02 03 04 05 06
SLB 1 [rA] = + 02 04 06 08 10 [rX] = - 12 14 16 18 20

2.1.2.11 Miscellaneous operators

Finally, we list in the following table three miscellaneous MIX instructions which do not fit in any of the previous subsections:

MOVE Move MOD words from M to the location stored in rI1. OPCODE = 7, MOD = no. of words.

NOP No operation. OPCODE = 0, MOD = 0.

HLT Halt. Stops instruction fetching. OPCODE = 5, MOD = 2.

The only effect of executing ‘NOP’ is increasing the location counter, while ‘HLT’ usually marks program termination.

2.1.2.12 Execution times

When writing MIXAL programs (or any kind of programs, for that matter), we shall often be interested in their execution time. Loosely speaking, we will be interested in the answer to the question: how long does it take a program to execute? Of course, this execution time will be a function of the input size, and the answer to our question is commonly given as the asymptotic behaviour as a function of the input size. At any rate, to compute this asymptotic behaviour, we need a measure of how long execution of a single instruction takes in our (virtual) CPU. Therefore, each MIX instruction will have an associated execution time, given in arbitrary units (in a real computer, the value of this unit will depend on the hardware configuration). When our MIX virtual machine executes programs, it will (optionally) give you the value of their execution time based upon the execution time of each single instruction.

In the following table, the execution times (in the above mentioned arbitrary units) of the MIX instructions are given.

NOP  1 ADD  2 SUB  2 MUL  10
DIV  12 NUM  10 CHAR  10 HLT  10
SLx  2 SRx  2 LDx  2 STx  2
JBUS  1 IOC  1 IN  1 OUT  1
JRED  1 Jx  1 INCx  1 DECx  1
ENTx  1 ENNx  1 CMPx  1 MOVE  1+2F

In the above table, ’F’ stands for the number of blocks to be moved (given by the FSPEC subfield of the instruction); SLx and SRx are a short cut for the byte-shifting operations; LDx denote all the loading operations; STx are the storing operations; Jx stands for all the jump operations, and so on with the rest of abbreviations.

2.2 MIXAL

In the previous sections we have listed all the available MIX binary instructions. As we have shown, each instruction is represented by a word which is fetched from memory and executed by the MIX virtual CPU. As is the case with real computers, the MIX knows how to decode instructions in binary format (the so–called machine language), but a human programmer would have a tough time if she were to write her programs in machine language. Fortunately, the MIX computer can be programmed using an assembly language, MIXAL, which provides a symbolic way of writing the binary instructions understood by the imaginary MIX computer. If you have used assembler languages before, you will find MIXAL a very familiar language. MIXAL source files are translated to machine language by a MIX assembler, which produces a binary file (the actual MIX program) which can be directly loaded into the MIX memory and subsequently executed.

In this section, we describe MIXAL, the MIX assembly language. The implementation of the MIX assembler program and MIX computer simulator provided by mdk are described later on (see Chapter 3 [Getting started], page 27).

2.2.1 Basic program structure

The MIX assembler reads MIXAL files line by line, producing, when required, a binary instruction, which is associated to a predefined memory address. To keep track of the current address, the assembler maintains an internal location counter which is incremented each time an instruction is compiled. In addition to MIX instructions, you can include in MIXAL file assembly directives (or pseudoinstructions) addressed at the assembler itself (for instance, telling it where the program starts and ends, or to reposition the location counter; see below).

MIX instructions and assembler directives^{4} are written in MIXAL (one per source file line) according to the following pattern:

     [LABEL]   MNEMONIC  [OPERAND]   [COMMENT]

where ‘OPERAND’ is of the form

     [ADDRESS][,INDEX][(MOD)]

Items between square brackets are optional, and

LABEL is an alphanumeric identifier (a symbol) which gets the current value of the location counter, and can be used in subsequent expressions,

MNEMONIC is a literal denoting the operation code of the instruction (e.g. LDA, STA; see see Section 2.1.2 [MIX instruction set], page 11) or an assembly pseudoinstruction (e.g. ORG, EQU),

ADDRESS is an expression evaluating to the address subfield of the instruction,

INDEX is an expression evaluating to the index subfield of the instruction, which defaults to 0 (i.e., no use of indexing) and can only be used when ADDRESS is present,

MOD is an expression evaluating to the mod subfield of the instruction. Its default value, when omitted, depends on OPCODE,

COMMENT any number of spaces after the operand mark the beginning of a comment, i.e. any text separated by white space from the operand is ignored by the assembler (note that spaces are not allowed within the ‘OPERAND’ field).

Note that spaces are not allowed between the ADDRESS, INDEX and MOD fields if they are present. White space is used to separate the label, operation code and operand parts of the instruction^{5}.

We have already listed the mnemonics associated with each MIX instruction; sample MIXAL instructions representing MIX instructions are:

     HERE     LDA  2000         HERE represents the current location counter
              LDX  HERE,2(1:3)  this is a comment
              JMP  1234

2.2.2 MIXAL directives

MIXAL instructions can be either one of the MIX machine instructions (see Section 2.1.2 [MIX instruction set], page 11) or one of the following assembly pseudoinstructions:

ORIG Sets the value of the memory address to which following instructions will be allocated after compilation.

EQU Used to define a symbol’s value, e.g. SYM EQU 2*200/3.

CON The value of the given expression is copied directly into the current memory address.

4 We shall call them, collectively, MIXAL instructions.

5 In fact, Knuth’s definition of MIXAL restricts the column number at which each of these instruction

parts must start. The MIXAL assembler included in mdk, mixasm, does not impose such restriction.

ALF Takes as operand five characters, constituting the five bytes of a word which is copied directly into the current memory address.

END Marks the end of the program. Its operand gives the start address for program execution.

The operand of ORIG, EQU, CON and END can be any expression evaluating to a constant MIX word, i.e., either a simple MIXAL expression (composed of numbers, symbols and binary operators, see Section 2.2.3 [Expressions], page 22) or a w-expression (see Section 2.2.4 [W-expressions], page 23).

All MIXAL programs must contain an END directive, with a twofold end: first, it marks the end of the assembler job, and, in the second place, its (mandatory) operand indicates the start address for the compiled program (that is, the address at which the virtual MIX machine must begin fetching instructions after loading the program). It is also very common (although not mandatory) to include at least an ORIG directive to mark the initial value of the assembler’s location counter (remember that it stores the address associated with each compiled MIX instruction). Thus, a minimal MIXAL program would be

               ORIG  2000    set the initial compilation address
               NOP           this instruction will be loaded at address 2000
               HLT           and this one at address 2001
               END   2000    end of program; start at address 2000
     this line is not parsed by the assembler

The assembler will generate two binary instructions (NOP (+ 00 00 00 00 00) and HLT (+ 00 00 02 05)), which will be loaded at addresses 2000 and 2001. Execution of the program will begin at address 2000. Every MIXAL program should also include a HLT instruction, which will mark the end of program execution (but not of program compilation).

The EQU directive allows the definition of symbolic names for specific values. For instance, we could rewrite the above program as follows:

     START     EQU   2000
               ORIG  START
               NOP
               HLT
               END   START

which would give rise to the same compiled code. Symbolic constants (or symbols, for short) can also be implicitly defined placing them in the LABEL field of a MIXAL instruction: in this case, the assembler assigns to the symbol the value of the location counter before compiling the line. Hence, a third way of writing our trivial program is

               ORIG  2000
     START     NOP
               HLT
               END   START

The CON directive allows you to directly specify the contents of the memory address pointed by the location counter. For instance, when the assembler encounters the following code snippet

               ORIG  1150
               CON   -1823473

it will assign to the memory cell number 1150 the contents - 00 06 61 11 49 (which corresponds to the decimal value -1823473).

Finally, the ALF directive lets you specify the memory contents as a set of five (optionally quoted) characters, which are translated by the assembler to their byte values, conforming in that way the binary word that is to be stored in the corresponding memory cell. This directive comes in handy when you need to store printable messages in a memory address, as in the following example^{6}:

               OUT  MSG       MSG is not yet defined here (future reference)
     MSG       ALF  "THIS "   MSG gets defined here
               ALF  "IS A "
               ALF  "MESSA"
               ALF  "GE.  "

The above snippet also shows the use of a future reference, that is, the usage of a symbol (MSG in the example) prior of its actual definition. The MIXAL assembler is able to handle future references subject to some limitations which are described in the following section (see Section 2.2.3 [Expressions], page 22).

Any line starting with an asterisk is treated as a comment and ignored by the assembler.

     * This is a comment: this line is ignored.
         * This line is an error: * must be in column 1.

As noted in the previous section, comments can also be located after the OPERAND field of an instruction, separated from it by white space, as in

     LABEL     LDA   100  This is also a comment

2.2.3 Expressions

The ADDRESS, INDEX and MOD fields of a MIXAL instruction can be expressions, formed by numbers, identifiers and binary operators (+ - * / // :). + and - can also be used as unary operators. Operator precedence is from left to right: there is no other operator precedence rule, and parentheses cannot be used for grouping. A stand-alone asterisk denotes the current memory location; thus, for instance,

          4+2**

evaluates to 6 (4 plus 2) times the current memory location. White space is not allowed within expressions.

The special binary operator : has the same meaning as in fspecs, i.e.,

     A:B = 8*A + B

while A//B stands for the quotient of the ten-byte number A 00 00 00 00 00 (that is, A right-padded with 5 null bytes or, what amounts to the same, multiplied by 64 to the fifth power) divided by B. Sample expressions are:

     18-8*3 = 30
     14/3 = 4
     1+3:11 = 4:11 = 43

6 In the original MIXAL definition, the ALF argument is not quoted. You can write the operand (as the

ADDRESS field) without quotes, but, in this case, you must follow the alignment rules of the original

MIXAL definition (namely, the ADDRESS must start at column 17).

     1//64 = (01 00 00 00 00 00)/(00 00 00 01 00) = (01 00 00 00 00)

Note that all MIXAL expressions evaluate to a MIX word (by definition).

All symbols appearing within an expression must be previously defined. Future references are only allowed when appearing standalone (or modified by an unary operator) in the ADDRESS part of a MIXAL instruction, e.g.

     * OK: stand alone future reference
              STA  -S1(1:5)
     * ERROR: future reference in expression
              LDX  2-S1
     S1       LD1  2000

2.2.4 W-expressions

Besides expressions, as described above (see Section 2.2.3 [Expressions], page 22), the MIXAL assembler is able to handle the so called w-expressions as the operands of the directives ORIG, EQU, CON and END (see Section 2.2.2 [MIXAL directives], page 20). The general form of a w-expression is the following:

          WEXP = EXP[(EXP)][,WEXP]

where EXP stands for an expression and square brackets denote optional items. Thus, a wexpression is made by an expression, followed by an optional expression between parenthesis, followed by any number of similar constructs separated by commas. Sample w-expressions are:

     2000
     235(3)
     S1+3(S2),3000
     S1,S2(3:5),23

W-expressions are evaluated from left to right as follows:

• Start with an accumulated result ‘w’ equal to 0.

• Take the first expression of the comma-separated list and evaluate it. For instance, if the w-expression is ‘S1+2(2:4),2000(S2)’, we evaluate first ‘S1+2’; let’s suppose that

    ‘S1’ equals 265230: then ‘S1+2 = 265232 = + 00 01 00 48 16’.

• Evaluate the expression within parenthesis, reducing it to an f-spec of the form ‘L:R’. In our previous example, the expression between parenthesis already has the desired form: 2:4.

• Substitute the bytes of the accumulated result ‘w’ designated by the f-spec using those of the previous expression value. In our sample, ‘w = + 00 00 00 00 00’, and we must substitute bytes 2, 3 and 4 of ‘w’ using values from 265232. We need 3 bytes, and we take the least significant ones: 00, 48, and 16, and insert them in positions 2, 3 and 4 of ‘w’, obtaining ‘w = + 00 00 48 16 00’.

• Repeat this operation with the remaining terms, acting on the new value of ‘w’. In our example, if, say, ‘S2 = 1:1’, we must substitute the first byte of ‘w’ using one byte (the least significant) from 2000, that is, 16 (since 2000 = + 00 00 00 31 16) and, therefore, we obtain ‘w = + 16 00 48 16 00’; summing up, we have obtained

    ‘265232(1:4),2000(1:1) = + 16 00 48 16 00 = 268633088’.

As a second example, in the w-expression

1(1:2),66(4:5)

we first take two bytes from 1 (00 and 01) and store them as bytes 1 and 2 of the result (obtaining ‘+ 00 01 00 00 00’) and, afterwards, take two bytes from 66 (01 and 02) and store them as bytes 4 and 5 of the result, obtaining ‘+ 00 01 00 01 02’ (262210). The process is repeated for each new comma-separated example. For instance:

1(1:1),2(2:2),3(3:3),4(4:4) = 01 02 03 04 00

As stated before, w-expressions can only appear as the operands of MIXAL directives taking a constant value (ORIG, EQU, CON and END). Future references are not allowed within w-expressions (i.e., all symbols appearing in a w-expression must be defined before it is used).

2.2.5 Local symbols

Besides user defined symbols, MIXAL programmers can use the so called local symbols, which are symbols of the form [1-9][HBF]. A local symbol nB refers to the address of the last previous occurrence of nH as a label, while nF refers to the next nH occurrence. Unlike user defined symbols, nH can appear multiple times in the LABEL part of different MIXAL instructions. The following code shows an instance of local symbols’ usage:

* line 1
1H    LDA  100
* line 2: 1B refers to address of line 1, 3F refers to address of line 4
      STA  3F,2(1B//2)
* line 3: redefinition of 1H
1H    STZ
* line 4: 1B refers to address of line 3
3H    JMP  1B

Note that a B local symbol never refers to a definition in its own line, that is, in the following program:

ORIG 1999
ST NOP
3H EQU 69
3H ENTA 3B  local symbol 3B refers to 3H in previous line
HLT
END ST

the contents of ‘rA’ is set to 69 and not to 2001. An specially tricky case occurs when using local symbols in conjunction with ORIG pseudoinstructions. To wit^{7},

ORIG 1999
ST NOP
3H CON 10
ENT1 *
LDA 3B
** rI1 is 2001, rA is 10.  So far so good!
3H ORIG 3B+1000

7 The author wants to thank Philip E. King for pointing these two special cases of local symbol usage to

him.

** at this point 3H equals 2003
** and the location counter equals 3000.
ENT2 *
LDX 3B
** rI2 contains 3000, rX contains 2003.
HLT
END ST

2.2.6 Literal constants

MIXAL allows the introduction of literal constants, which are automatically stored in memory addresses after the end of the program by the assembler. Literal constants are denoted as =wexp=, where wexp is a w-expression (see Section 2.2.4 [W-expressions], page 23). For instance, the code

L         EQU   5
          LDA   =20-L=

causes the assembler to add after the program’s end an instruction with contents 15 (‘20-L’), and to assemble the above code as the instruction LDA a, where a stands for the address in which the value 15 is stored. In other words, the compiled code is equivalent to the following:

L         EQU  5
          LDA  a
...
a         CON  20-L
          END  start

3 Getting started

In this chapter, you will find a sample code-compile-run-debug session using the mdk utilities. Familiarity with the MIX mythical computer and its assembly language MIXAL (as described in Knuth’s TAOCP) is assumed; for a compact reminder, see Chapter 2 [MIX and MIXAL tutorial], page 9.

3.1 Writing a source file

MIXAL programs can be written as ASCII files with your editor of choice. Here you have the mandatory hello world as written in the MIXAL assembly language:

     *                                                        (1)
     * hello.mixal: say 'hello world' in MIXAL                (2)
     *                                                        (3)
     * label ins    operand     comment                       (4)
     TERM    EQU    19          the MIX console device number (5)
             ORIG   3000        start address                 (6)
     START   OUT    MSG(TERM)   output data at address MSG    (7)
             HLT                halt execution                (8)
     MSG     ALF    "MIXAL"                                   (9)
             ALF    " HELL"                                   (10)
             ALF    "O WOR"                                   (11)
             ALF    "LD   "                                   (12)
             END    START       end of the program            (13)

MIXAL source files should have the extension .mixal when used with the mdk utilities. As you can see in the above sample, each line in a MIXAL file can be divided into four fields separated by an arbitrary amount of whitespace characters (blanks and or tabs). While in Knuth’s definition of MIXAL each field must start at a fixed pre-defined column number, the mdk assembler loosens this requirement and lets you format the file as you see fit. The only restrictions retained are for comment lines (like 1-4) which must begin with an asterisk (*) placed at column 1, and for the label field (see below) which, if present, must also start at column 1. The four fields in each non-comment line are:

− an optional label, which either refers to the current memory address (as START and MSG in lines 7 and 9) or a defined symbol (TERM) (if present, the label must always start at the first column in its line, for the first whitespace in the line marks the beginning of the second field),

− an operation mnemonic, which can represent either a MIX instruction (OUT and HLT in lines 7 and 8 above), or an assembly pseudoinstruction (e.g., the ORIG pseudoinstruction in line 6^{1}.

− an optional operand for the (pseudo)instruction, and

− an optional free text comment.

1 If an ORIG directive is not used, the program will be loaded by the virtual machine at address 0. ORIG

allows allocating the executable code where you see fit.

Lines 9-12 of the hello.mixal file above also show the second (and last) difference between Knuth’s MIXAL definition and ours: the operand of the ALF pseudoinstruction (a word of five characters) must be quoted using ""^{2}.

The workings of this sample program should be straightforward if you are familiar with MIXAL. See TAOCP vol. 1 for a thorough definition or Chapter 2 [MIX and MIXAL tutorial], page 9, for a tutorial.

3.2 Compiling

Three simulators of the MIX computer, called mixvm, gmixvm and mixguile, are included in the mdk tools. They are able to run binary files containing MIX instructions written in their binary representation. You can translate MIXAL source files into this binary form using mixasm, the MIXAL assembler. So, in order to compile the hello.mixal file, you can type the following command at your shell prompt:

     mixasm hello RET

If the source file contains no errors, this will produce a binary file called hello.mix which can be loaded and run by the MIX virtual machine. Unless the mixasm option -O is provided, the assembler will include debug information in the executable file (for a complete description of all the compilation options, see Chapter 5 [mixasm], page 45). Now, your are ready to run your first MIX program, as described in the following section.

3.3 Running the program

MIX is a mythical computer, so it is no use ordering it from your favorite hardware provider. mdk provides three software simulators of the computer, though. They are

• mixvm, a command line oriented simulator,

• gmixvm, a GTK based graphical interface to mixvm, and

• mixguile, a Guile shell with a built-in MIX simulator.

All three simulators accept the same set of user commands, but offer a different user interface, as noted above. In this section we shall describe some of these commands, and show you how to use them from mixvm’s command line. You can use them as well at gmixvm’s command prompt (see Chapter 7 [gmixvm], page 59), or using the built-in Scheme primitives of mixguile (see Section 3.4 [Using mixguile], page 33).

Using the MIX simulators, you can run your MIXAL programs, after compiling them with mixasm into binary .mix files. mixvm can be used either in interactive or non-interactive mode. In the second case, mixvm will load your program into memory, execute it (producing any output due to MIXAL OUT instructions present in the program), and exit when it encounters a HLT instruction. In interactive mode, you will enter a shell prompt which allows you issuing commands to the running virtual machine. These commands will permit you to load, run and debug programs, as well as to inspect the MIX computer state (register contents, memory cells contents and so on).

2 In Knuth’s definition, the operand always starts at a fixed column number, and the use of quotation

is therefore unnecessary. As mixasm releases this requirement, marking the beginning and end of the

ALF operand disambiguates the parser’s recognition of this operand when it includes blanks. Note that

double-quotes (") are not part of the MIX character set, and, therefore, no escape characters are needed

within ALF’s operands.

3.3.1 Non-interactive mode

To make mixvm work in non-interactive mode, use the -r flag. Thus, to run our hello.mix program, simply type

     mixvm -r hello RET

at your command prompt, and you will get the following output:

     MIXAL HELLO WORLD

Since our hello world program uses MIX’s device number 19 as its output device (see Section 3.1 [Writing a source file], page 27), the output is redirected to the shell’s standard output. Had you used any other MIX output devices (disks, drums, line printer, etc.), mixvm would have created a file named after the device used (e.g. disk4.dev) and written its output there^{3}.

The virtual machine can also report the execution time of the program, according to the (virtual) time spent in each of the binary instructions (see Section 2.1.2.12 [Execution times], page 18). Printing of execution time statistics is activated with the -t flag; running

     mixvm -t -r hello RET

produces the following output:

     MIXAL HELLO WORLD
     ** Execution time: 11

Sometimes, you will prefer to store the results of your program in MIX registers rather than writing them to a device. In such cases, mixvm’s -d flag is your friend: it makes mixvm dump the contents of its registers and flags after executing the loaded program. For instance, typing the following command at your shell’s prompt

     mixvm -d -r hello

you will obtain the following output:

     MIXAL HELLO WORLD
     rA: + 00 00 00 00 00 (0000000000)
     rX: + 00 00 00 00 00 (0000000000)
     rJ: + 00 00 (0000)
     rI1: + 00 00 (0000)     rI2: + 00 00 (0000)
     rI3: + 00 00 (0000)     rI4: + 00 00 (0000)
     rI5: + 00 00 (0000)     rI6: + 00 00 (0000)
     Overflow: F
     Cmp: E

which, in addition to the program’s outputs and execution time, gives you the contents of the MIX registers and the values of the overflow toggle and comparison flag (admittedly, rather uninteresting in our sample).

As you can see, running programs non-interactively has many limitations. You cannot peek the virtual machine’s memory contents, not to mention stepping through your program’s instructions or setting breakpoints^{4}. Enter interactive mode.

3 The device files are stored, by default, in a directory called .mdk, which is created in your home directory

the first time mixvm is run. You can change this default directory using the command devdir when

running mixvm in interactive mode (see Section 6.2.4 [Configuration commands], page 56)

4 The mixguile program allows you to execute arbitrary combinations of mixvm commands (using Scheme)

non-interactively. See Section 3.4.5 [Scheme scripts], page 39.

3.3.2 Interactive mode

To enter the MIX virtual machine interactive mode, simply type

     mixvm RET

at your shell command prompt. This command enters the mixvm command shell. You will be presented the following command prompt:

     MIX >

The virtual machine is initialised and ready to accept your commands. The mixvm command shell uses GNU’s readline, so that you have at your disposal command completion (using TAB) and history functionality, as well as other line editing shortcuts common to all utilities using this library (for a complete description of readline’s line editing usage, see Section “Command Line Editing” in Readline.)

Usually, the first thing you will want to do is loading a compiled MIX program into memory. This is accomplished by the load command, which takes as an argument the name of the .mix file to be loaded. Thus, typing

     MIX > load hello RET
     Program loaded. Start address: 3000
     MIX >

will load hello.mix into the virtual machine’s memory and set the program counter to the address of the first instruction. You can obtain the contents of the program counter using the command pc:

     MIX > pc
     Current address: 3000
     MIX >

After loading it, you are ready to run the program, using, as you surely have guessed, the run command:

     MIX > run
     Running ...
     MIXAL HELLO WORLD
     ... done
     Elapsed time: 11 /Total program time: 11 (Total uptime: 11)
     MIX >

Note that now the timing statistics are richer. You obtain the elapsed execution time (i.e., the time spent executing instructions since the last breakpoint), the total execution time for the program up to now (which in our case coincides with the elapsed time, since there were no breakpoints), and the total uptime for the virtual machine (you can load and run more than one program in the same session)^{5}. After running the program, the program counter will point to the address after the one containing the HLT instruction. In our case, asking the value of the program counter after executing the program will give us

     MIX > pc
     Current address: 3002
     MIX >

5 Printing of timing statistics can be disabled using the command timing (see Section 6.2.4 [Configuration

commands], page 56).

You can check the contents of a memory cell giving its address as an argument of the command pmem, like this

MIX > pmem 3001
3001: + 00 00 00 02 05 (0000000133)
MIX >

and convince yourself that address 3001 contains the binary representation of the instruction HLT. An address range of the form FROM-TO can also be used as the argument of pmem:

MIX > pmem 3000-3006
3000: + 46 58 00 19 37 (0786957541)
3001: + 00 00 00 02 05 (0000000133)
3002: + 14 09 27 01 13 (0237350989)
3003: + 00 08 05 13 13 (0002118477)
3004: + 16 00 26 16 19 (0268542995)
3005: + 13 04 00 00 00 (0219152384)
3006: + 00 00 00 00 00 (0000000000)
MIX >

In a similar manner, you can look at the contents of the MIX registers and flags. For instance, to ask for the contents of the A register you can type

MIX > preg A
rA: + 00 00 00 00 00 (0000000000)
MIX >

Use the command help to obtain a list of all available commands, and help COMMAND for help on a specific command, e.g.

MIX > help run
run             Run loaded or given MIX code file. Usage: run [FILENAME]
MIX >

For a complete list of commands available at the MIX propmt, See Chapter 6 [mixvm], page 47. In the following subsection, you will find a quick tour over commands useful for debugging your programs.

3.3.3 Debugging commands

The interactive mode of mixvm lets you step by step execution of programs as well as breakpoint setting. Use next to step through the program, running its instructions one by one. To run our two-instruction hello.mix sample you can do the following:

MIX > load hello
Program loaded. Start address: 3000
MIX > pc
Current address: 3000
MIX > next
MIXAL HELLO WORLD
Elapsed time: 1 /Total program time: 1 (Total uptime: 1)
MIX > pc
Current address: 3001
MIX > next
End of program reached at address 3002

Elapsed time: 10 /Total program time: 11 (Total uptime: 11)
MIX > pc
Current address: 3002
MIX > next
MIXAL HELLO WORLD
Elapsed time: 1 /Total program time: 1 (Total uptime: 12)
MIX >
MIX > run
Running ...
... done
Elapsed time: 10 /Total program time: 11 (Total uptime: 22)
MIX >

(As an aside, the above sample also shows how the virtual machine handles cumulative time statistics and automatic program restart).

You can set a breakpoint at a given address using the command sbpa (set breakpoint at address). When a breakpoint is set, run will stop before executing the instruction at the given address. Typing run again will resume program execution. Coming back to our hello world example, we would have:

MIX > sbpa 3001
Breakpoint set at address 3001
MIX > run
Running ...
MIXAL HELLO WORLD
... stopped: breakpoint at line 8 (address 3001)
Elapsed time: 1 /Total program time: 1 (Total uptime: 23)
MIX > run
Running ...
... done
Elapsed time: 10 /Total program time: 11 (Total uptime: 33)
MIX >

Note that, since we compiled hello.mixal with debug info enabled, the virtual machine is able to tell us the line in the source file corresponding to the breakpoint we are setting. As a matter of fact, you can directly set breakpoints at source code lines using the command

sbp LINE_NO, e.g.
MIX > sbp 4
Breakpoint set at line 7
MIX >

sbp sets the breakpoint at the first meaningful source code line; thus, in the above example we have requested a breakpoint at a line which does not correspond to a MIX instruction and the breakpoint is set at the first line containing a real instruction after the given one. To unset breakpoints, use cbpa ADDRESS and cbp LINE_NO, or cabp to remove all currently set breakpoints. You can also set conditional breakpoints, i.e., tell mixvm to interrupt program execution whenever a register, a memory cell, the comparison flag or the overflow toggle change using the commands sbp[rmco] (see Section 6.2.2 [Debug commands], page 50).

MIXAL lets you define symbolic constants, either using the EQU pseudoinstruction or starting an instruction line with a label (which assigns to the label the value of the current memory address). Each MIXAL program has, therefore, an associated symbol table which you can inspect using the psym command. For our hello world sample, you will obtain the following output:

     MIX > psym
     START:  3000
     TERM:  19
     MSG:  3002
     MIX >

Other useful commands for debugging are strace (which turns on tracing of executed instructions), pbt (which prints a backtrace of executed instructions) and weval (which evaluates w-expressions on the fly). For a complete description of all available MIX commands, See Chapter 6 [mixvm], page 47.

3.4 Using mixguile

With mixguile you can run a MIX simulator embedded in a Guile shell, that is, using Scheme functions and programs. As with mixvm, mixguile can be run both in interactive and non-interactive modes. The following subsections provide a quick tour on using this MIX emulator.

3.4.1 The mixguile shell

If you simply type

     mixguile RET

at the command prompt, you’ll be presented a Guile shell prompt like this

     guile>

At this point, you have entered a Scheme read-eval-print loop (REPL) which offers you all the Guile functionality plus a new set of built-in procedures to execute and debug MIX programs. Each of the mixvm commands described in the previous sections (and in see Chapter 6 [mixvm], page 47) have a Scheme function counterpart named after it by prepending the prefix mix- to its name. Thus, to load our hello world program, you can simply enter

     guile> (mix-load "hello")
     Program loaded. Start address: 3000
     guile>

and run it using mix-run:

     guile> (mix-run)
     Running ...
     MIXAL HELLO WORLD
     ... done
     Elapsed time: 11 /Total program time: 11 (Total uptime: 11)
     guile>

In the same way, you can execute it step by step using the Scheme function mix-next or set a breakpoint:

     guile> (mix-sbp 4)
     Breakpoint set at line 5
     guile>

or, if you one to peek at a register contents:

     guile> (mix-preg 'A)
     rA: + 00 00 00 00 00 (0000000000)
     guile>

You get the idea: you have at your disposal all the mixvm and gmixvm commands by means of mix- functions. But, in case you are wondering, this is only the beginning. You also have at your disposal a whole Scheme interpreter, and you can, for instance, define new functions combining the mix- and all other Scheme primitives. In the next sections, you’ll find examples of how to take advantage of the Guile interpreter.

3.4.2 Additional MIX Scheme functions

The mix- function counterparts of the mixvm commands don’t return any value, and are evaluated only for their side-effects (possibly including informational messages to the standard output and/or error stream). When writing your own Scheme functions to manipulate the MIX virtual machine within mixguile (see Section 3.4.3 [Defining new functions], page 35), you’ll probably need Scheme functions returning the value of the registers, memory cells and so on. Don’t worry: mixguile also offers you such functions. For instance, to access the (numerical) value of a register you can use mix-reg:

     guile> (mix-reg 'I2)
     0
     guile>

Note that, unlike (mix-preg 'I2), the expression (mix-reg 'I2) in the above example evaluates to a Scheme number and does not produce any side-effect:

     guile> (number? (mix-reg 'I2))
     #t
     guile> (number? (mix-preg 'I2))
     rI2: + 00 00 (0000)
     #f
     guile>

In a similar fashion, you can access the memory contents using (mix-cell), or the program counter using (mix-loc):

     guile> (mix-cell 3000)
     786957541
     guile> (mix-loc)
     3002
     guile>

Other functions returning the contents of the virtual machine components are mix-cmp and mix-over, which eval to the value of the comparison flag and the overflow toggle respectively. For a complete list of these additional functions, See Chapter 8 [mixguile], page 65.

In the next section, we’ll see a sample of using these functions to extend mixguile’s functionality.

3.4.3 Defining new functions

Scheme is a powerful language, and you can use it inside mixguile to easily extend the MIX interpreter’s capabilities. For example, you can easily define a function that loads a file, prints its name, executes it and, finally, shows the registers contents, all in one shot:

guile> (define my-load-and-run  RET
         (lambda (file)   RET
           (mix-load file)   RET
           (display "File loaded: ")   RET
           (mix-pprog)   RET
           (mix-run)   RET
           (mix-preg)))   RET
guile>

and use it to run your programs:

guile> (my-load-and-run "hello")
Program loaded. Start address: 3000
File loaded: hello.mix
Running ...
MIXAL HELLO WORLD
... done
Elapsed time: 11 /Total program time: 11 (Total uptime: 33)
rA: + 00 00 00 00 00 (0000000000)
rX: + 00 00 00 00 00 (0000000000)
rJ: + 00 00 (0000)
rI1: + 00 00 (0000) rI2: + 00 00 (0000)
rI3: + 00 00 (0000) rI4: + 00 00 (0000)
rI5: + 00 00 (0000) rI6: + 00 00 (0000)
guile>

Or, maybe, you want a function which sets a breakpoint at a specified line number before executing it:

guile> (define my-load-and-run-with-bp
         (lambda (file line)
           (mix-load file)
           (mix-sbp line)
           (mix-run)))
guile> (my-load-and-run-with-bp "samples/primes" 10)
Program loaded. Start address: 3000
Breakpoint set at line 10
Running ...
... stopped: breakpoint at line 10 (address 3001)
Elapsed time: 1 /Total program time: 1 (Total uptime: 45)
guile>

As a third example, the following function loads a program, runs it and prints the contents of the memory between the program’s start and end addresses:

guile> (define my-run
         (lambda (file)

                (mix-load file)
                (let ((start (mix-loc)))
                  (mix-run)
                  (mix-pmem start (mix-loc)))))
     guile> (my-run "hello")
     Program loaded. Start address: 3000
     Running ...
     MIXAL HELLO WORLD
     ... done
     Elapsed time: 11 /Total program time: 11 (Total uptime: 11)
     3000: + 46 58 00 19 37 (0786957541)
     3001: + 00 00 00 02 05 (0000000133)
     3002: + 14 09 27 01 13 (0237350989)
     guile>

As you can see, the possibilities are virtually unlimited. Of course, you don’t need to type a function definition each time you start mixguile. You can write it in a file, and load it using Scheme’s load function. For instance, you can create a file named, say, functions.scm with your definitions (or any Scheme expression) and load it at the

mixguile prompt:
     guile> (load "functions.scm")

Alternatively, you can make mixguile to load it for you. When mixguile starts, it looks for a file named mixguile.scm in your MDK configuration directory (~/.mdk) and, if it exists, loads it before entering the REPL. Therefore, you can copy your definitions in

that file, or load the functions.scm file in mixguile.scm.

3.4.4 Hook functions

Hooks are functions called before or after a given event occurs. In mixguile, you can define command and break hooks, which are associated, respectively, with command execution and program interruption events. The following sections give you a tutorial on using hook functions within mixguile.

3.4.4.1 Command hooks

In the previous section, we have seen how to extend mixguile’s functionality through the use of user defined functions. Frequently, you’ll write new functions that improve in some way the workings of a built-in mixvm command, following this pattern:

a. Prepare the command execution b. Execute the desired command c. Perform post execution operations

We call the functions executed in step (a) pre-hooks, and those of step post-hooks of the given command. mixguile lets you specify pre- and post-hooks for any mixvm command using the mix-add-pre-hook and mix-add-post-hook functions, which take as arguments a symbol naming the command and a function to be executed before (resp. after) the command. In other words, mixguile will execute for you steps (a) and (c) above whenever you eval (b). The hook functions must take a single argument, which is a string list of the command’s arguments. As an example, let us define the following hooks for the next command:

     (define next-pre-hook
       (lambda (arglist)
         (mix-slog #f)))
     (define next-post-hook
       (lambda (arglist)
         (display "Stopped at line ")
         (display (mix-src-line-no))
         (display ": ")
         (display (mix-src-line))
         (newline)
         (mix-slog #t)))

In these functions, we are using the function mix-slog to turn off the informational messages produced by the virtual machine, since we are providing our own ones in the post hook function. To install these hooks, we would write:

     (mix-add-pre-hook 'next next-pre-hook)
     (mix-add-post-hook 'next next-post-hook)

Assuming we have put the above expressions in mixguile’s initialisation file, we would obtain the following results when evaluating mix-next:

     guile> (mix-next)
     MIXAL HELLO WORLD
     Stopped at line 6:             HLT
     guile>

As a second, more elaborate, example, let’s define hooks which print the address and contents of a cell being modified using smem. The hook functions could be something like this:

     (define smem-pre-hook
       (lambda (arglist)
         (if (eq? (length arglist) 2)
             (begin
               (display "Changing address ")
               (display (car arglist))
               (newline)
               (display "Old contents: ")
               (display (mix-cell (string->number (car arglist))))
               (newline))
             (error "Wrong arguments" arglist))))
     (define smem-post-hook
       (lambda (arglist)
         (if (eq? (length arglist) 2)
             (begin
               (display "New contents: ")

          (display (mix-cell (string->number (car arglist))))
          (newline)))))

and we can install them using

(mix-add-pre-hook 'smem smem-pre-hook)
(mix-add-post-hook 'smem smem-post-hook)

Afterwards, a sample execution of mix-smem would look like this:

guile> (mix-smem 2000 100)
Changing address 2000
Old contents: 0
New contents: 100
guile>

You can add any number of hooks to a given command. They will be executed in the same order as they are registered. You can also define global post (pre) hooks, which will be called before (after) any mixvm command is executed. Global hook functions must admit two arguments, namely, a string naming the invoked command and a string list of its arguments, and they are installed using the Scheme functions mix-add-global-pre-hook and mix-add-global-post-hook. A simple example of global hook would be:

guile> (define pre-hook
         (lambda (cmd args)
           (display cmd)
           (display " invoked with arguments ")
           (display args)
           (newline)))
guile> (mix-add-global-pre-hook pre-hook)
ok
guile> (mix-pmem 120 125)
pmem invoked with arguments (120-125)
0120: + 00 00 00 00 00 (0000000000)
0121: + 00 00 00 00 00 (0000000000)
0122: + 00 00 00 00 00 (0000000000)
0123: + 00 00 00 00 00 (0000000000)
0124: + 00 00 00 00 00 (0000000000)
0125: + 00 00 00 00 00 (0000000000)
guile>

Note that if you invoke mixvm commands within a global hook, its associated command hooks will be run. Thus, if you have installed both the next hooks described earlier and the global hook above, executing mix-next will yield the following result:

guile> (mix-next 5)
next invoked with arguments (5)
slog invoked with arguments (off)
MIXAL HELLO WORLD
Stopped at line 7: MSG         ALF   "MIXAL"
slog invoked with arguments (on)
guile>

Adventurous readers may see the above global hook as the beginning of a command log utility or a macro recorder that saves your commands for replay.

3.4.4.2 Break hooks

We have seen in the previous section how to associate hooks to command execution, but they are not the whole story. You can also associate hook functions to program interruption, that is, specify functions that should be called every time the execution of a MIX program is stopped due to the presence of a breakpoint, either explicit or conditional. Break hooks take as arguments the line number and memory address at which the break occurred. A simple hook that logs the line and address of the breakpoint could be defined as:

     (define break-hook
       (lambda (line address)
         (display "Breakpoint encountered at line ")
         (display line)
         (display " and address ")
         (display address)
         (newline)))

and installed for explicit and conditional breakpoints using

     (mix-add-break-hook break-hook)
     (mix-add-cond-break-hook break-hook)

after that, every time the virtual machine encounters a breakpoint, break-code shall be evaluated for you^{6}.

3.4.5 Scheme scripts

Another useful way of using mixguile is writing executable scripts that perform a set of commands for you. This is done using the mixguile switch -s (being a Guile shell, mixguile accepts all the command options of guile; type mixguile -h for a list of all available command options). For instance, if you have a very useful MIX program foo.mix which you want to run often, you don’t have to fire up a MIX virtual machine, load and run it every time; you can write a Scheme script instead:

     #! /usr/bin/mixguile -s
     !#
     ;;; runprimes: execute the primes.mix program
     ;; load the file you want to run
     (mix-load "../samples/primes")
     ;; execute it
     (mix-run)
     ;; print the contents of registers
     (mix-pall)
     ;; ...

6 You may have noticed that break hooks can be implemented in terms of command hooks associated to

mix-run and mix-next. As a matter of fact, they are implemented this way: take a look at the file

install_dir/share/mdk/mix-vm-stat.scm if you are curious.

Just save the above script to a file named, say, runtest, make it executable (chmod +x runtest), and, well, execute it from the Unix shell:

$ ./runtest
Program loaded. Start address: 3000
Running ...
... done
Elapsed time: 190908 /Total program time: 190908 (Total uptime: 190908)
rA: + 30 30 30 30 30 (0511305630)
rX: + 30 30 32 32 39 (0511313959)
rJ: + 47 18 (3026)
rI1: + 00 00 (0000)     rI2: + 55 51 (3571)
rI3: + 00 19 (0019)     rI4: + 31 51 (2035)
rI5: + 00 00 (0000)     rI6: + 00 00 (0000)
Overflow: F
Cmp: L
$

Note that this is far more flexible that running programs non-interactively using mixvm (see Section 3.3.1 [Non-interactive mode], page 29), for you can execute any combination of commands you want from a Scheme script (not just running and dumping the registers). For additional mixguile command line options, see Section 8.1 [Invoking mixguile], page 65.

3.5 Using Scheme in mixvm and gmixvm

In the previous section (see Section 3.4 [Using mixguile], page 33) we have seen how the Guile shell mixguile offers you the possibility of using Scheme to manipulate a MIX virtual machine and extend the set of commands offered by mixvm and gmixvm. This possibility is not limited to the mixguile shell. Actually, both mixvm and gmixvm incorporate an embedded Guile interpreter, and can evaluate Scheme expressions. To evaluate a single-line expression at the mixvm or gmixvm command prompt, simply write it and press return (the command parser will recognise it as a Scheme expression because it is parenthesized, and will pass it to the Guile interpreter). A sample mixvm session using Scheme expressions could be:

MIX > load hello
Program loaded. Start address: 3000
MIX > (define a (mix-loc))
MIX > run
Running ...
MIXAL HELLO WORLD
... done
Elapsed time: 11 /Total program time: 11 (Total uptime: 11)
MIX > (mix-pmem a)
3000: + 46 58 00 19 37 (0786957541)
MIX > (mix-pmem (mix-loc))
3002: + 14 09 27 01 13 (0237350989)
MIX >

You can also load and evaluate a file, using the scmf command like this:

     MIX> scmf /path/to/file/file.scm

Therefore, you have at your disposal all the mixguile goodies described above (new functions, new command definitions, hooks...) inside mixvm and gmixvm. In other words, these programs are extensible using Scheme. See Section 3.4 [Using mixguile], page 33 for examples of how to do it.

4 Emacs tools

Everyone writing code knows how important a good editor is. Most systems already come with Emacs, and excellent programmer’s editor. mdk adds support to Emacs for both writing and debugging MIX programs. A major mode for MIXAL source files eases edition of your code, while integration with Emacs’ debugging interface (GUD) lets you use mixvm without leaving your favourite text editor.

This chapter shows how to use the Elisp modules included in mdk, assuming that you have followed the installation instructions in See Section 1.4 [Emacs support], page 6.

4.1 MIXAL mode

The module mixal-mode.el provides a new mode, mixal-mode, for editing MIXAL source files^{1}. When everything is installed correctly, Emacs will select it as the major mode for editing files with extension .mixal. You can also activate mixal-mode in any buffer issuing the Emacs command M-x mixal-mode.

4.1.1 Basics

The mode for editing mixal source files is inherited from fundamental-mode, meaning that all your favorite editing operations will still work. If you want a short introduction to Emacs, type C-h t inside Emacs to start the tutorial.

Mixal mode adds font locking. If you do not have font locking globally enabled, you can turn it on for mixal-mode by placing the following line in your .emacs file:

     (add-hook 'mixal-mode-hook 'turn-on-font-lock)

You can also customize the colors used to colour your mixal code by changing the requisite faces. This is the list of faces used by mixal-mode:

• font-lock-comment-face Face to use for comments.

• mixal-font-lock-label-face Face to use for label names.

• mixal-font-lock-operation-code-face Face to use for operation code names.

• mixal-font-lock-assembly-pseudoinstruction-face Face to use for assembly pseudo-instruction names.

4.1.2 Help system

When coding your program, you will be thinking, looking up documentation and editing files. Emacs already helps you with editing files, but Emacs can do much more. In particular, looking up documentation is one of its strong points. Besides the info system (which you are probably already using), mixal-mode defines commands for getting particular information about a MIX operation code.

   With M-x mixal-describe-operation-code (or its keyboard shortcut C-h o) you will

get the documentation about a particular MIX operation code. Keep in mind that these are not assembly (MIXAL) pseudoinstructions. When the point is around a MIXAL pseudoinstruction in your source file, Emacs will recognize it and will suggest the right MIX operation code.

1 mixal-mode has been developed and documented by Pieter E. J. Pareit

4.1.3 Compiling and running

After you have written your MIXAL program, you’ll probably want to test it. This can be done with the MIX virtual machine. First you will need to compile your code into MIX byte code. This can be done within Emacs with the command M-x compile (C-c c). In case of compilation errors, you can jump to the offending source code line with M-x next-error.

Once the program compiles without errors, you can debug or run it. To invoke the debugger, use M-x mixal-debug (C-c d). Emacs will open a GUD buffer where you can use the debugging commands described in See Chapter 6 [mixvm], page 47.

If you just want to execute the program, you can do so with M-x mixal-run (C-c r). This will invoke mixvm, execute the program and show its output in a separate buffer.

4.2 GUD integration

If you are an Emacs user and write your MIXAL programs using this editor, you will find the elisp program mixvm.el quite useful^{2}. mixvm.el allows running the MIX virtual machine mixvm (see Chapter 6 [mixvm], page 47) inside an Emacs GUD buffer, while visiting the MIXAL source file in another buffer.

After installing mixvm.el (see Section 1.4 [Emacs support], page 6), you can initiate an mdk/GUD session inside Emacs with the command

     M-x mixvm

and you will have a mixvm prompt inside a newly created GUD buffer. GUD will reflect the current line in the corresponding source file buffer.

2 mixvm.el has been kindly contributed by Philip E. King. mixvm.el is based on a study of gdb, perldb,

and pdb as found in gud.el, and rubydb3x.el distributed with the source code to the Ruby language.

5 mixasm, the MIXAL assembler

MIX programs, as executed by mixvm, are composed of binary instructions loaded into the virtual machine memory as MIX words. Although you could write your MIX programs directly as a series of words in binary format, you have at your disposal a more friendly assembly language, MIXAL (see Section 2.2 [MIXAL], page 19) which is compiled into binary form by mixasm, the MIXAL assembler included in mdk. In this chapter, you will find a complete description of mixasm options.

5.1 Invoking mixasm

In its simplest form, mixasm is invoked with a single argument, which is the name of the MIXAL file to be compiled, e.g.

     mixasm hello

will compile either hello or hello.mixal, producing a binary file named hello.mix if no errors are found.

In addition, mixasm can be invoked with the following command line options (note, that, following GNU’s conventions, we provide a long option name for each available single letter switch):

     mixasm [-vhulO] [-o OUTPUT_FILE] [--version] [--help] [--usage]
            [--ndebug] [--output=OUTPUT_FILE] [--list[=LIST_FILE]] file

The meaning of these options is as follows:

[User Option] -v [User Option] --version

Prints version and copyleft information and exits.

[User Option] -h [User Option] --help [User Option] -u [User Option] --usage

Prints a summary of available options and exits.

[User Option] -O [User Option] --ndebug

Do not include debugging information in the compiled file, saving space but disallowing breakpoint setting at source level and symbol table inspection under mixvm.

[User Option] -o output file

[User Option] --output=output_file

By default, the given source file file.mixal is compiled into file.mix. You can provide a different name for the output file using this option.

[User Option] -l

--list[=list_file]  [User Option]

This option causes mixasm to produce, in addition to the .mix file, an ASCII file containing a summary of the compilation results. The file is named after the MIXAL source file, changing its extension to .mls if no argument is provided; otherwise, the listing file is named according to the argument.

6 mixvm, the MIX computer simulator

This chapter describes mixvm, the MIX computer simulator. mixvm is a command line interface programme which simulates the MIX computer (see Section 2.1 [The MIX computer], page 9). It is able to run MIXAL programs (see Section 2.2 [MIXAL], page 19) previously compiled with the MIX assembler (see Chapter 5 [mixasm], page 45). The simulator allows inspection of the MIX computer components (registers, memory cells, comparison flag and overflow toggle), step by step execution of MIX programmes, and breakpoint setting to aid you in debugging your code. For a tutorial description of mixvm usage, See Section 3.3 [Running the program], page 28.

6.1 Invoking mixvm

mixvm can be invoked with the following command line options (note that, following GNU’s conventions, we provide a long option name for each available single letter switch):

mixvm [-vhurdtq] [--version] [--help] [--usage] [--run] [--dump]
      [--time] [--noinit]  [FILE[.mix]]

The meaning of these options is as follows:

[User Option] -v [User Option] --version

Prints version and copyleft information and exits.

[User Option] -h [User Option] --help [User Option] -u [User Option] --usage

Prints a summary of available options and exits.

[User Option] -r [User Option] --run

Loads the specified FILE and executes it. After the program execution, mixvm exits. FILE must be the name of a binary .mix program compiled with mixasm. If your program does not produce any output, use the -d flag (see below) to peek at the virtual machine’s state after execution.

[User Option] -d [User Option] --dump

This option must be used in conjunction with -r, and tells mixvm to print the value of the virtual machine’s registers, comparison flag and overflow toggle after executing the program named FILE. See See Section 3.3.1 [Non-interactive mode], page 29, for sample usage.

[User Option] -t [User Option] --time

This option must be used in conjunction with -r, and tells mixvm to print virtual time statistics for the program’s execution.

When run without the -r flag, mixvm enters its interactive mode, showing you a prompt like this one:

     MIX >

and waiting for your commands (see Section 6.2 [Commands], page 48). If the optional FILE argument is given, the file FILE.mix will be loaded into the virtual machine memory before entering the interactive mode.

The first time mixvm is invoked, a directory named .mdk is created in your home directory. It contains the mixvm configuration file, the command history file and (by default) the block devices files (see Section 6.3 [Devices], page 57). Before showing you the command prompt, mixvm looks in the ~/.mdk directory for a file named mixguile.scm; if it exists, it is read and evaluated by the embedded Guile interpreter (see Section 3.4.3 [Defining new functions], page 35). You can use the -q command line option to skip this file loading:

[User Option] -q [User Option] --noinit

Do not load the Guile initialisation file ~/.mdk/mixguile.scm at startup.

6.2 Interactive commands

You can enter the interactive mode of the MIX virtual machine by simply invoking mixvm without arguments. You will then be greeted by a shell prompt^{1}

     MIX >

which indicates that a new virtual machine has been initialised and is ready to execute your commands. As we have already mentioned, this command prompt offers you command line editing facilities which are described in the Readline user’s manual (chances are that you are already familiar with these command line editing capabilities, as they are present in many GNU utilities, e.g. the bash shell)^{2}. In a nutshell, readline provides command completion using the TAB key and command history using the cursor keys. A history file containing the last commands typed in previous sessions is stored in the mdk configuration directory

(~/.mdk).

As a beginner, your best friend will be the help command, which shows you a summary of all available MIX commands and their usage; its syntax is as follows:

[mixvm command] help [command] Prints a short description of the given command and its usage. If command is omitted, help prints the short description for all available commands.

6.2.1 File commands

You have at your disposal a series of commands that let you load and execute MIX executable files, as well as manipulate MIXAL source files:

1 The default command prompt, ‘MIX > ’, can be changed using the prompt command (see Section 6.2.4

[Configuration commands], page 56)

2 The readline functionality will be available if you have compiled mdk with readline support, i.e., if GNU

readline is installed in your system. This is often the case in GNU/Linux and BSD systems

[file command] load file[.mix] This command loads a binary file, file.mix into the virtual machine memory, and positions the program counter at the beginning of the loaded program. This address is indicated in the MIXAL source file as the operand of the END pseudoinstruction. Thus, if your sample.mixal source file contains the line:

          END 3000

and you compile it with mixasm to produce the binary file sample.mix, you will load it into the virtual machine as follows:

     MIX > load sample
     Program loaded. Start address: 3000
     MIX >

[file command] run [file[.mix]]

When executed without argument, this command initiates or resumes execution of instructions from the current program counter address. Therefore, issuing this command after a successful load, will run the loaded program until either a HLT instruction or a breakpoint is found. If you provide a MIX filename as argument, the given file will be loaded (as with load file) and executed. If run is invoked again after program execution completion (i.e., after the HLT instruction has been found in a previous run), the program counter is repositioned and execution starts again from the beginning (as a matter of fact, a load command preserving the currently set breakpoints is issued before resuming execution).

[file command] edit [file[.mixal]] The source file file.mixal is edited using the editor defined in the environment variable MDK EDITOR. If this variable is not set, the following ones are tried out in order: X EDITOR, EDITOR and VISUAL. If invoked without argument, the source file for the currently loaded MIX file is edited. The command used to edit source files can also be configured using the sedit command (see Section 6.2.4 [Configuration commands], page 56).

[file command] compile file[.mixal]

The source file file.mixal is compiled (with debug information enabled) using mixasm. If invoked without argument, the source file for the currently loaded MIX file is recompiled. The compilation command can be set using the sasm command (see Section 6.2.4 [Configuration commands], page 56).

[file command] pprog [file command] psrc

Print the path of the currently loaded MIX program and its source file:

     MIX > load ../samples/primes
     Program loaded. Start address: 3000
     MIX > pprog
     ../samples/primes.mix
     MIX > psrc
     /home/jao/projects/mdk/gnu/samples/primes.mixal
     MIx>

Finally, you can use the quit command to exit mixvm:

[file command] quit

Exit mixvm, saving the current configuration parameters in ~/.mdk/mixvm.config.

6.2.2 Debug commands

Sequential execution of loaded programs can be interrupted using the following debug commands:

[debug command] next [ins number] This command causes the virtual machine to fetch and execute up to ins number instructions, beginning from the current program counter position. Execution is interrupted either when the specified number of instructions have been fetched or a breakpoint is found, whatever happens first. If run without arguments, one instruction is executed. If next is invoked again after program execution completion (i.e., after the HLT instruction has been found in a previous run), the program counter is repositioned and execution starts again from the beginning (as a matter of fact, a load command preserving the currently set breakpoints is issued before resuming execution).

[debug command] sbp line number [debug command] cbp line no

Sets a breakpoint at the specified source file line number. If the line specified corresponds to a command or to a MIXAL pseudoinstruction which does not produce a MIX instruction in the binary file (such as ORIG or EQU) the breakpoint is set at the first source code line giving rise to a MIX instruction after the specified one. Thus, for our sample hello.mixal file:

     *                                                        (1)
     * hello.mixal: say 'hello world' in MIXAL                (2)
     *                                                        (3)
     * label ins    operand     comment                       (4)
     TERM    EQU    19          the MIX console device number (5)
             ORIG   1000        start address                 (6)
     START   OUT    MSG(TERM)   output data at address MSG    (7)
     ...

trying to set a breakpoint at line 5, will produce the following result:

     MIX > sbp 5
     Breakpoint set at line 7
     MIX >

since line 7 is the first one compiled into a MIX instruction (at address 3000). The command cbp clears a (previously set) breakpoint at the given source file line.

[debug command] spba address [debug command] cbpa address

Sets a breakpoint at the given memory address. The argument must be a valid MIX memory address, i.e., it must belong into the range [0-3999]. Note that no check is performed to verify that the specified address is reachable during program execution.

No debug information is needed to set a breakpoint by address with sbpa. The command cbpa clears a (previously set) breakpoint at the given memory address.

[debug command] sbpr A | X | J | Ii [debug command] cbpr A | X | J | Ii

Sets a conditional breakpoint on the specified register change. For instance,

     sbpr I1

will cause an interruption during program execution whenever the contents of register I1 changes. A previously set breakpoint is cleared using the cbpr command.

[debug command] sbpm address [debug command] cbpm address

Sets a conditional breakpoint on the specified memory cell change. The argument must be a valid MIX memory address, i.e., it must belong into the range [0-3999]. For instance,

     sbpm 1000

will cause an interruption during program execution whenever the contents of the memory cell number 1000 changes. A previously set breakpoint is cleared using the cbpm command.

[debug command] sbpo [debug command] cbpo

Sets/clears a conditional breakpoint on overflow toggle change.

[debug command] sbpc [debug command] cbpc

Sets/clears a conditional breakpoint on comparison flag change.

[debug command] cabp

Clears all currently set breakpoints.

[debug command] psym [symbol name] MIXAL programs can define symbolic constants, using either the EQU pseudoinstruction or a label at the beginning of a line. Thus, in the program fragment

     VAR     EQU  2168
             ORIG 4000
     START   LDA  VAR

the symbol VAR stands for the value 2168, while START is assigned the value 4000. The symbol table can be consulted from the mixvm command line using psym followed by the name of the symbol whose contents you are interested in. When run without arguments, psym will print all defined symbols and their values.

The virtual machine can also show you the instructions it is executing, using the following commands:

[debug command] strace [on|off]

strace on enables instruction tracing. When tracing is enabled, each time the virtual machine executes an instruction (due to your issuing a run or next command), it is printed in its canonical form (that is, with all expressions evaluated to their numerical values) and, if the program was compiled with debug information, as it was originally typed in the MIXAL source file. Instruction tracing is disabled with strace off command. A typical tracing session could be like this:

MIX > strace on
MIX > next
3000: [OUT 3002,0(2:3)] START OUT MSG(TERM)
MIXAL HELLO WORLD
Elapsed time: 1 /Total program time: 1 (Total uptime: 1)
MIX > next
3001: [HLT 0,0] HLT
End of program reached at address 3002
Elapsed time: 10 /Total program time: 11 (Total uptime: 11)
MIX > strace off
MIX >

The executed instruction, as it was translated, is shown between square brackets after the memory address, and, following it, you can see the actual MIXAL code that was compiled into the executed instruction. The tracing behaviour is stored as a configuration parameter in ~/.mdk.

[debug command] pline [LINE NUMBER]

Prints the requested source line (or the current one if line number is omitted:

MIX > load ../samples/hello
Program loaded. Start address: 3000
MIX > pline
Line 5: START       OUT   MSG(TERM)
MIX > pline 6
Line 6:             HLT
MIX >

[debug command] pbt [INS NUMBER] This command prints a backtrace of executed instructions. Its optional argument ins number is the number of instructions to print. If it is omitted or equals zero, all executed instructions are printed. For instance, if you compile and load the following program (bt.mixal):

    ORIG 0
BEG JMP *+1
    JMP *+1
FOO JMP BAR
BAR HLT
    END BEG

you could get the following traces:

MIX > load bt
Program loaded. Start address: 0
MIX > next
MIX > pbt

#0      BEG     in bt.mixal:2
MIX > next
MIX > pbt
#0      1       in bt.mixal:3
#1      BEG     in bt.mixal:2
MIX > run
Running ...
... done
MIX > pbt 3
#0      BAR     in bt.mixal:5
#1      FOO     in bt.mixal:4
#2      1       in bt.mixal:3
MIX > pbt
#0      BAR     in bt.mixal:5
#1      FOO     in bt.mixal:4
#2      1       in bt.mixal:3
#3      BEG     in bt.mixal:2
MIX >

Note that the executed instruction trace gives you the label of the executed line or, if it has no label, its address.

As you have probably observed, mixvm prints timing statistics when running programs. This behaviour can be controlled using the stime command (see Section 6.2.4 [Configuration commands], page 56).

mixvm is also able of evaluating w-expressions (see Section 2.2.4 [W-expressions], page 23) using the following command:

[debug command] weval WEXP

Evaluates the given w-expression, WEXP. The w-expression can contain any currently defined symbol. For instance:

MIX > psym START
+ 00 00 00 46 56 (0000003000)
MIX > weval START(0:1),START(3:4)
+ 56 00 46 56 00 (0939716096)
MIX >

New symbols can be defined using the ssym command:

[debug command] ssym SYM WEXP Defines the symbol named SYM with the value resulting from evaluating WEXP, a w-expression. The newly defined symbol can be used in subsequent weval commands, as part of the expression to be evaluated. E.g.,

MIX > ssym S 2+23*START
+ 00 00 18 19 56 (0000075000)
MIX > psym S
+ 00 00 18 19 56 (0000075000)
MIX > weval S(3:4)

+ 00 00 19 56 00 (0000081408)
MIX >

Finally, if you want to discover which is the decimal value of a MIX word expressed as five bytes plus sign, you can use

[debug command] w2d WORD Computes the decimal value of the given word. WORD must be expressed as a sign (+/-) followed by five space-delimited, two-digit decimal values representing the five bytes composing the word. The reverse operation (showing the word representation of a decimal value) can be accomplished with weval. For instance:

MIX > w2d - 01 00 00 02 02
-16777346
MIX > weval -16777346
- 01 00 00 02 02 (0016777346)
MIX >

6.2.3 State commands

Inspection and modification of the virtual machine state (memory, registers, overflow toggle and comparison flag contents) is accomplished using the following commands:

[state command] pstat

This commands prints the current virtual machine state, which can be one of the following: − No program loaded − Program successfully loaded − Execution stopped (next executed) − Execution stopped: breakpoint encountered − Execution stopped: conditional breakpoint encountered − Program successfully terminated

[state command] pc

Prints the current value of the program counter, which stores the address of the next instruction to be executed in a non-halted program.

[state command] sreg A | X | J | I[1-6] value

[state command] preg [A | X | J | I[1-6]] [state command] pall

preg prints the contents of a given MIX register. For instance, preg A will print the contents of the A-register. When invoked without arguments, all registers shall be printed:

MIX > preg
rA: - 00 00 00 00 35 (0000000035)
rX: + 00 00 00 15 40 (0000001000)
rJ: + 00 00 (0000)
rI1: + 00 00 (0000) rI2: + 00 00 (0000)
rI3: + 00 00 (0000) rI4: + 00 00 (0000)

     rI5: + 00 00 (0000) rI6: + 00 00 (0000)
     MIX >

As you can see in the above sample, the contents are printed as the sign plus the values of the MIX bytes stored in the register and, between parenthesis, the decimal representation of its module.

pall prints the contents of all registers plus the comparison flag and overflow toggle.

Finally, sreg Sets the contents of the given register to value, expressed as a decimal constant. If value exceeds the maximum value storable in the given register, VALUE

mod MAXIMUM_VALUE is stored, e.g.
     MIX > sreg I1 1000
     MIX > preg I1
     rI1: + 15 40 (1000)
     MIX > sreg I1 1000000
     MIX > preg I1
     rI1: + 09 00 (0576)
     MIX >

[state command] pflags [state command] scmp E | G | L [state command] sover F | T

pflags prints the value of the comparison flag and overflow toggle of the virtual machine, e.g.

     MIX > pflags
     Overflow: F
     Cmp: E
     MIX >

The values of the overflow toggle are either F (false) or T (true), and, for the comparison flag, E, G, L (equal, greater, lesser). scmp and sover are setters of the comparison flag and overflow toggle values.

[state command] pmem from[-to] [state command] smem address value

pmem prints the contents of memory cells in the address range [FROM-TO]. If the upper limit to is omitted, only the contents of the memory cell with address FROM is printed, as in

     MIX > pmem 3000
     3000: + 46 58 00 19 37 (0786957541)
     MIX >

The memory contents are displayed both as the set of five MIX bytes plus sign composing the stored MIX word and, between parenthesis, the decimal representation of the module of the stored value.

smem sets the content of the memory cell with address address to value, expressed as a decimal constant.

6.2.4 Configuration commands

This section describes commands that allow you to configure the virtual machine behaviour. This configuration is stored in the mdk directory ~/.mdk.

As you can see in their description, some commands print, as a side effect, informational messages to the standard output (e.g. load prints a message telling you the loaded program’s start address): these messages can be enabled/disabled using slog:

[config command] slog on|off

Turns on/off the logging of informational messages. Note that error messages are always displayed, as well as state messages required using commands prefixed with p (preg, pmem and the like).

[config command] stime on|off [config command] ptime

The stime command (un)sets the printing of timing statistics, and ptime prints their current value:

     MIX > ptime
     Elapsed time: 10 /Total program time: 11 (Total uptime: 11)
     MIX >

[config command] sedit TEMPLATE

[config command] pedit

sedit sets the command to be used to edit MIXAL source files with the edit command. TEMPLATE must contain the control characters %s to mark the place where the source’s file name will be inserted. For instance, if you type

     MIX > sedit emacsclient %s
     MIX >

issuing the mixvm command edit foo.mixal will invoke the operating system com-

mand emacsclient foo.mixal.

pedit prints the current value of the edit command template.

[config command] sasm TEMPLATE

[config command] pasm

sasm sets the command to be used to compile MIXAL source files with the compile command. template must contain the control characters %s to mark the place where the source’s file name will be inserted. For instance, if you type

     MIX > sasm mixasm -l %s
     MIX >

issuing the mixvm command compile foo.mixal will invoke the operating system

command mixasm -l foo.mixal.

pasm prints the current value of the compile command template.

[config command] sddir DIRNAME

[config command] pddir

MIX devices (see Section 6.3 [Devices], page 57) are implemented as regular files stored, by default, inside ~/.mdk. The sddir command lets you specify an alternative location for storing these device files, while pddir prints the current device directory.

Finally, you can change the default command prompt, ‘MIX > ’, using the prompt command:

[config command] prompt PROMPT

Changes the command prompt to prompt. If you want to include white space(s) at the end of the new prompt, bracket prompt using double quotes (e.g., prompt ">> ").

6.2.5 Scheme commands

If you have compiled mdk with libguile support (see Section 1.5 [Special configure flags], page 7), mixvm will start and initialise an embedded Guile Scheme interpreter when it is invoked. That means that you have at your disposal, at mixvm’s command prompt, all the Scheme primitives described in Section 3.4 [Using mixguile], page 33 and Chapter 8 [mixguile], page 65, as well as any other function or hook that you have defined in the initialisation file ~/.mdk/mixguile.scm. To evaluate a Scheme function, simply type it at the mixvm command prompt (see Section 3.5 [Using Scheme in mixvm and gmixvm], page 40 for a sample). Compared to the mixguile program, this has only one limitation: the expressions used in mixvm cannot span more than one line. You can get over this inconvenience writing your multiline Scheme expressions in a file and loading it using the scmf command:

[scheme command] scmf FILE NAME Loads the given Scheme file and evaluates it using the embedded Guile interpreter.

6.3 MIX block devices

The MIX computer comes equipped with a set of block devices for input-output operations (see Section 2.1.2.8 [Input-output operators], page 16). mixvm implements these block devices as disk files, with the exception of block device no. 19 (typewriter terminal) which is redirected to standard input/output. When you request an output operation on any other (output) device, a file named according to the following table will be created, and the specified MIX words will be written to the file in binary form (for binary devices) or in ASCII (for char devices). Files corresponding to input block devices should be created and filled beforehand to be used by the MIX virtual machine (for input-output devices this creation can be accomplished by a MIXAL program writing to the device the required data, or, if you prefer, with your favourite editor). The device files are stored, by default, in the directory ~/.mdk; this location can be changed using the mixvm command devdir (see Section 6.2.4 [Configuration commands], page 56).

Device  No.  filename  type  and block size
Tape  0-7  tape[0-7].dev    bin i/o - 100 words
Disks  8-15  disk[0-7].dev    bin i/o - 100 words
Card reader  16  cardrd.dev       char in - 16 words
Card writer  17  cardwr.dev       char out - 16 words

Line printer  18  printer.dev      char out - 24 words
Terminal  19  stdin/stdout     char i/o - 14 words
Paper tape  20  paper.dev        char in - 14 words

Devices of type char are stored as ASCII files, using one line per block. For instance, since the card reader has blocks of size 16, that is, 80 characters, it will be emulated by an ASCII file consisting of lines with length 80. If the reader finds a line with less than the required number of characters, it pads the memory with zeroes (MIX character ’space’) to complete the block size.

Note that the virtual machine automatically converts between the MIX and ASCII character encodings, so that you can manipulate char device files with any ASCII editor. In addition, the reader is not case-sensitive, i.e., it automatically converts lowercase letters to their uppercase counterparts (since the MIX character set does not include the former).

The typewriter (device no. 19) lets you use the standard input and output in your MIXAL programs. For instance, here is a simple ’echo’ program:

     * simple echo program
     TERM    EQU     19          the typewriter device
     BUF     EQU     500         input buffer
             ORIG    1000
     START   IN      BUF(TERM)   read a block (70 chars)
             OUT     BUF(TERM)   write the read chars
             HLT
             END     START

Input lines longer than 70 characters (14 words) are trimmed. On the other hand, if you type less than a block of characters, whitespace (MIX character zero) is used as padding.

7 gmixvm, the GTK virtual machine

This chapter describes the graphical MIX virtual machine emulator shipped with mdk. In addition to having all the command-oriented functionalities of the other virtual machines (mixvm and mixguile), gmixvm offers you a graphical interface displaying the status of the virtual machine, the source code of the the downloaded programs and the contents of the MIX devices.

7.1 Invoking gmixvm

If you have built mdk with GTK+ support (see Chapter 1 [Installing MDK], page 5), a graphical front-end for the MIX virtual machine will be available in your system. You can invoke it by typing

     gmixvm [-vhuq] [--version] [--help] [--usage] [--noinit]

at your command prompt, where the options have the following meanings:

[User Option] -v [User Option] --version

Prints version and copyleft information and exits.

[User Option] -h [User Option] --help [User Option] -u [User Option] --usage

Prints a summary of available options and exits.

[User Option] -q [User Option] --noinit

Do not load the Guile initialisation file ~/.mdk/mixguile.scm at startup. This file contains any local Scheme code to be executed by the embedded Guile interpreter at startup (see Section 3.5 [Using Scheme in mixvm and gmixvm], page 40).

Typing gmixvm or gmixvm -q at your command prompt, the main window will appear, offering you a graphical interface to run and debug your MIX programs.

Apart from the menu and status bars, we can distinguish two zones (or halves) in this main window. In the upper half of gmixvm’s main window there is a notebook with three pages, namely,

• a MIX virtual machine view, which shows you the registers, flags, memory contents and time statistics of the virtual machine;

• a MIXAL source view, which shows the MIXAL file and lets you manage breakpoints;

• a Devices view, which shows you the output to character based MIX block devices.

These three windows can be detached from the notebook, using either the penultimate toolbar button (which detachs the currently visible notebook page) or the menu entries

under View->Detached windows.

On the other hand, the main window’s lower half presents you a mixvm command prompt and a logging area where results of the issued commands are presented. These widgets implement a mixvm console which offers almost the same functionality as its CLI counterpart.

When gmixvm is run, it creates a directory named .mdk in your home directory (if it does not already exist). The .mdk directory contains the program settings, the device files used by your MIX programs (see Section 6.3 [Devices], page 57), and a command history file.

The following sections describe the above mentioned components of gmixvm.

7.2 MIXVM console

In the lower half of the gmixvm main window, you will find a command text entry and, above it, an echo area. These widgets offer you the same functionality as its CLI counterpart, mixvm (see Chapter 6 [mixvm], page 47). You can issue almost all mixmv commands at the gmixvm’s command prompt in order to manipulate the MIX virtual machine. Please refer to See Chapter 6 [mixvm], page 47, for a description of these commands, and to See Chapter 3 [Getting started], page 27, for a tutorial on using the MIX virtual machine. The command prompt offers command line completion for partially typed commands using the TAB key; e.g., if you type

     lo TAB

the command is automatically completed to load. If multiple completions are available, they will be shown in the echo area. Thus, typing

     p TAB

will produce the following output on the echo area:

     Completions:
     pc     psym     preg     pflags     pall
     pmem

which lists all the available commands starting with p. In addition, the command prompt maintains a history of typed commands, which can be recovered using the arrow up and down keys. As mentioned above, a file containing previous sessions’ commands is stored in the configuration directory ~/.mdk, and reloaded every time you start gmixvm.

You can change the font used to display the issued commands and the mes-

sages in the echo area using the Settings->Change font->Command prompt and
Settings->Change font->Command log menu commands.

7.3 MIX virtual machine

The first notebook’s page displays the current status of the virtual machine. There you can find the registers’ contents, the value of the comparison and overflow flags, the location pointer, a list with all MIX memory cells and their contents, and the time statistics (including total uptime, elapsed time since the last run command and total execution time for the currently loaded MIX program).

If you click any register entry, you will be prompted for a new register’s contents.

In the same manner, click on any address of the memory cells list to be prompted for the new contents of the clicked cell. If you click the address column’s title, a dialog asking you for a memory address will appear; if you introduce a valid address, this will be the first cell displayed in the scrollable list after you click the OK button.

The register contents are shown as a list of MIX bytes plus sign. If you place the mouse pointer over any of them, the decimal value of this MIX word will appear inside a tooltip.

You can change the font used to display the MIX virtual machine contents using the

Settings->Change font->MIX menu command.

7.4 MIXAL source view

The second notebook’s page, dubbed Source, shows you the MIXAL source of the currently loaded MIX file.

The information is presented in four columns. The first column displays little icons showing the current program pointer and any set breakpoints. The second and third columns show the address and memory contents of the compiled MIX instruction, while the last one displays its corresponding MIXAL representation, together with the source file line number. You can set/unset breakpoints by clicking on any line that has an associated memory address.

You can change the font used to display the MIXAL source code using the

Settings->Change font->MIXAL menu command.

7.5 MIX devices view

The last notebook page, dubbed Devices, shows you the output/input to/from MIX block devices (the console, line printer, paper tape, disks, card and tapes see Section 6.3 [Devices], page 57) produced by the running program.

Input device contents is read from files located in the ~/.mdk directory, and the output is also written to files at the same location. Note that device tabs will appear as they are used by the MIX program being run, and that loading a new MIX program will close all previously open devices.

The input/output for binary block devices (tapes and disks) is a list of MIX words, which can be displayed either in decimal or word format (e.g. - 67 or - 00 00 00 01 03). The format used by gmixvm can be configured using the Settings->Device output menu command for each binary device.

You can change the font used to display the devices content using the

Settings->Change font->Devices menu command.

7.6 Menu and status bars

The menu bar gives you access to the following commands:

[File] Load...

Opens a file dialog that lets you specify a binary MIX file to be loaded in the virtual machine’s memory. It is equivalent to the mixvm’s load command (see Section 6.2.1 [File commands], page 48).

[File] Edit...

Opens a file dialog that lets your specify a MIXAL source file to be edited. It is equivalent to the mixvm’s edit command (see Section 6.2.1 [File commands], page 48). The program used for editing can be specified using the menu entry

     Settings->External programs, or using the mixvm command sedit.

[File] Compile...

Opens a file dialog that lets your specify a MIXAL source file to be compiled. It is equivalent to the mixvm’s compile command (see Section 6.2.1 [File commands], page 48). The command used for compiling can be specified using the menu entry

Settings->External programs, or using the mixvm command sasm.

[File] Exit

Exits the application.

[Debug] Run

Runs the currently loaded MIX program, up to the next breakpoint. It is equivalent to the mixvm’s run command (see Section 6.2.2 [Debug commands], page 50).

[Debug] Next

Executes the next MIX instruction. It is equivalent to the mixvm’s next command (see Section 6.2.2 [Debug commands], page 50).

Clear breakpoints  [Debug]

Clears all currently set breakpoints. It is equivalent to the mixvm’s cabp command.

[Debug] Symbols...

Opens a dialog showing the list of symbols defined in the currently loaded MIX program. The font used to display this list can be customised using the meny entry

Settings->Change font->Symbol list.
[View] Toolbar(s)

Toggles the toolbar(s) in the gmixvm window(s) (when notebook pages are detached, each one has its own toolbar).

Detached windows Virtual machine  [View]

[View] Detached windows Source
[View] Detached windows Devices

These toggles let you detach (or re-attach) the corresponding notebook page.

[Settings] Change font

Lets you change the font used in the various gmixv widgets (i.e. commad prompt, command log, Virtual machine, Source, Devices and Symbol list). There is also an entry (All) to change all fonts at once.

Device output...  [Settings]

Opens a dialog that lets you specify which format shall be used to show the contents of MIX binary block devices. The available formats are decimal (e.g. -1234) and MIX word (e.g. - 00 00 00 19 18).

Devices dir...  [Settings]

Opens a dialog that lets you choose where the MIX device files will be stored (~/.mdk is the default location). You can also specify the devices directory using the mixvm command sddir (see Section 6.2.4 [Configuration commands], page 56).

External programs...  [Settings]

This menu command opens a dialog that lets you specify the commands used for editing and compiling MIXAL source files. The commands are specified as template strings, where the control substring %s will be substituted by the actual file name. Thus, if you want to edit programs using vi running in an xterm, you must enter the command template xterm -e vi %s in the corresponding dialog entry. These settings can also be changed using the mixvm commands sedit and sasm (see Section 6.2.4 [Configuration commands], page 56).

[Settings] Save

Saves the current settings.

Save on exit  [Settings]

Mark this checkbox if you want gmixvm to save its settings every time you quit the program.

[Help] About...

Shows information about gmixvm’s version and copyright.

On the other hand, the status bar displays the name of the last loaded MIX file. In addition, when the mouse pointer is over a MIXAL source file line that contains symbols, a list of these symbols with their values will appear in the status bar.

8 mixguile, the Scheme virtual machine

This chapter provides a reference to using mixguile and the Scheme function library giving access to the MIX virtual machine in the mdk emulators (mixguile, mixvm and gmixvm). See Section 3.4 [Using mixguile], page 33 for a tutorial, step by step introduction to mixguile and using Scheme as an extension language for the mdk MIX virtual machines.

8.1 Invoking mixguile

Invoking mixguile without arguments will enter the Guile REPL (read-eval-print loop) after loading, if it exists, the user’s initialisation file (~/.mdk/mixguile.scm).

mixguile accepts the same command line options than Guile:

     mixguile [-s SCRIPT] [-c EXPR] [-l FILE] [-e FUNCTION] [-qhv]
              [--help] [--version]

The meaning of these options is as follows:

[User Option] -h [User Option] --help

Prints usage summary and exits.

[User Option] -v [User Option] --version

Prints version and copyleft information and exits.

[User Option] -s SCRIPT Loads Scheme code from script, evaluates it and exits. This option can be used to write executable Scheme scripts, as described in Section 3.4.5 [Scheme scripts], page 39.

[User Option] -c EXPR Evaluates the given Scheme expression and exits.

[User Option] -l FILE Loads the given Scheme file and enters the REPL (read-eval-print loop).

[User Option] -e FUNCTION After reading the script, executes the given function using the provided command line arguments. For instance, you can write the following Scheme script:

          #! /usr/bin/mixguile \
          -e main -s
          !#
          ;;; execute a given program and print the registers.
          (define main
            (lambda (args)
              ;; load the file provided as a command line argument
              (mix-load (cadr args))

              ;; execute it
              (mix-run)
              ;; print the contents of registers
              (mix-pall)))

save it in a file called, say, foo, make it executable, and run it as

          $ ./foo hello

This invocation will cause the evaluation of the main function with a list of command line parameters as its argument (("./foo" "hello") in the above example. Note that command line options to mixguile must be written in their own line after the \ symbol.

[User Option] -q

Do not load user’s initialisation file. When mixguile starts up, it looks for a file named mixguile.scm in the user’s mdk configuration directory (~/.mdk), and loads it if it exists. This option tells mixguile to skip this initialisation file loading.

8.2 Scheme functions reference

As we have previously pointed out, mixguile embeds a MIX virtual machine that can be accessed through a set of Scheme functions, that is, of a Scheme library. Conversely, mixvm and gmixvm contain a Guile interpreter, and are able to use this same Scheme library, as well as all the other Guile/Scheme primitives and any user defined function. Therefore, you have at your disposal a powerful programming language, Scheme, to extend the mdk virtual machine emulators (see Section 3.5 [Using Scheme in mixvm and gmixvm], page 40 for samples of how to do it).

The following subsections describe available functions the MIX/Scheme library.

8.2.1 mixvm command wrappers

For each of the mixvm commands listed in Section 6.2 [Commands], page 48, there is a corresponding Scheme function named by prefixing the command name with mix- (e.g., mix-load, mix-run and so on). These command wrappers are implemented using a generic command dispatching function:

[Function] mixvm-cmd command argument Dispatches the given command to the MIX virtual machine appending the provided argument. Both command and argument must be strings. The net result is as writing "command argument" at the mixvm or gmixvm command prompt.

For instance, you can invoke the run command at the mixvm prompt in three equivalent ways:

     MIX > run hello
     MIX > (mix-run "hello")
     MIX > (mixvm-cmd "run" "hello")

(only the two last forms can be used at the mixguile prompt or inside a Scheme script).

The mix- functions evaluate to a unspecified value. If you want to check the result of the last mixvm command invocation, use the mix-last-result function:

[Function] mix-last-result

Returns #t if the last mixvm command invocation was successful, #f otherwise.

Using this function, we could improve the script for running a program presented in the previous section by adding error checking:

#! /usr/bin/mixguile \
-e main -s
!#
;;; Execute a given program and print the registers.
(define main
  (lambda (args)
    ;; load the file provided as a command line argument
    (mix-load (cadr args))
    ;; execute it if mix-load succeeded
    (if (mix-last-result) (mix-run))
    ;; print the contents of registers if the above commands succeeded
    (if (mix-last-result) (mix-pall))))

Please, refer to Section 6.2 [Commands], page 48 for a list of available commands. Given the description of a mixvm, it is straightforward to use its Scheme counterpart and, therefore, we shall not give a complete description of these functions here. Instead, we will only mention those wrappers that exhibit a treatment of their differing from that of their command counterpart.

[Function] mix-preg [register]

[Function] mix-sreg register value The argument register of these functions can be either a string or a symbol representing the desired register. For instance, the following invocations are equivalent:

     (mix-preg 'I1)
     (mix-preg "I1")

[Function] mix-pmem from [to]

The command pmem takes a single argument which can be either a cell number or a range of the form FROM-TO. This function takes one argument to ask for a single memory cell contents, or two parameters to ask for a range. For instance, the following commands are equivalent:

     MIX > pmem 10-12
     0010: + 00 00 00 00 00 (0000000000)
     0011: + 00 00 00 00 00 (0000000000)
     0012: + 00 00 00 00 00 (0000000000)
     MIX > (mix-pmem 10 12)
     0010: + 00 00 00 00 00 (0000000000)
     0011: + 00 00 00 00 00 (0000000000)
     0012: + 00 00 00 00 00 (0000000000)

[Function] mix-sover #t|#f

The command sover takes as argument either the string T or the string F, to set, respectively, the overflow toggle to true or false. Its Scheme counterpart, mix-sover, takes as argument a Scheme boolean value: #t (true) or #f.

For the remaining functions, you simply must take into account that when the command arguments are numerical, the corresponding Scheme function takes as arguments Scheme number literals. On the other hand, when the command argument is a string, the argument of its associated Scheme function will be a Scheme string. By way of example, the following invocations are pairwise equivalent:

MIX > load ../samples/hello
MIX > (mix-load "../samples/hello")
MIX > next 5
MIX > (mix-next 5)

8.2.2 Hook functions

Hooks are functions evaluated before or after executing a mixvm command (or its corresponding Scheme function wrapper), or after an explicit or conditional breakpoint is found during the execution of a MIX program. The following functions let you install hooks:

[Function] mix-add-pre-hook command hook

Adds a function to the list of pre-hooks associated with the given command. command is a string naming the corresponding mixvm command, and hook is a function which takes a single argument: a string list of the commands arguments. The following scheme code defines a simple hook and associates it with the run command:

     (define run-hook
       (lambda (args)
         (display "argument list: ")
         (display args)
         (newline)))
     (mix-add-pre-hook "run" run-hook)

Pre-hooks are executed, in the order they are added, before invoking the corresponding command (or its associated Scheme wrapper function).

[Function] mix-add-post-hook command hook

Adds a function to the list of pre-hooks associated with the given command. The arguments have the same meaning as in mix-add-pre-hook.

[Function] mix-add-global-pre-hook hook
[Function] mix-add-global-post-hook hook

Global pre/post hooks are executed before/after any mixvm command or function wrapper invocation. In this case, hook takes two arguments: a string with the name of the command being invoked, and a string list with its arguments.

[Function] mix-add-break-hook hook
[Function] mix-add-cond-break hook

Add a hook funtion to be executed when an explicit (resp. conditional) breakpoint is encountered during program execution. hook is a function taking two arguments: the source line number where the hook has occurred, and the current program counter value. The following code shows a simple definition and installation of a break hook:

     (define break-hook
       (lambda (line address)
         (display "Breakpoint at line ") (display line)
         (display " and address ") (display address)
         (newline)))
     (mix-add-break-hook break-hook)

Break hook functions are entirely implemented in Scheme using regular post-hooks for the next and run commands. If you are curious, you can check the Scheme source

code at prefix/share/mdk/mixguile-vm-stat.scm (where prefix stands for your

root install directory, usually /usr or /usr/local.

See Section 3.4.4 [Hook functions], page 36 for further examples on using hook functions.

8.2.3 Additional VM functions

When writing non-trivial Scheme extensions using the MIX/Scheme library, you will probably need to evaluate the contents of the virtual machine components (registers, memory cells and so on). For instance, you may need to store the contents of the A register in a variable. The Scheme functions described so far are of no help: you can print the contents of A using (mix-preg 'A), but you cannot define a variable containing the contents of A. To address this kind of problems, the MIX/Scheme library provides the following additional functions:

[Function] mixvm-status
[Function] mix-vm-status

Return the current status of the virtual machine, as a number (mixvm-status) or as a symbol (mix-vm-status). Possible return values are:

(mixvm-status)     (mix-vm-status)

0  MIX ERROR  Loading or execution error
1  MIX BREAK  Breakpoint encountered
2  MIX COND BREAK  Conditional breakpoint
3  MIX HALTED  Execution terminated
4  MIX RUNNING  Execution stopped after next
5  MIX LOADED  Program successfully loaded
6  MIX EMPTY  No program loaded

[Function] mix-vm-error?
[Function] mix-vm-break?
[Function] mix-vm-cond-break?
[Function] mix-vm-halted?
[Function] mix-vm-running?
[Function] mix-vm-loaded?
[Function] mix-vm-empty?

Predicates asking whether the current virtual machine status is MIX_ERROR, MIX_

BREAK, etc.

[Function] mix-reg register

[Function] mix-set-reg! register value mix-reg evaluates to a number which is the contents of the specified register. mixset-reg sets the contents of the given register to value. The register can be specified either as a string ("A", "X", etc.) or as a symbol ('A, 'X, etc.). For instance,

     guile> (mix-reg 'A)
     2341
     guile> (mix-set-reg! "A" 2000)
     ok
     guile> (define reg-a (mix-reg 'A))
     guile> (display reg-a)
     2000
     guile>

[Function] mix-cell cell no

[Function] mix-set-cell! cell no value

Evaluate and set the contents of the memory cell number cell no. Both cell no and value are Scheme numbers.

[Function] mix-loc

Evaluates to the value of the location counter (i.e., the address of the next instruction to be executed).

[Function] mix-over

[Function] mix-set-over! #t|#f

mix-over evaluates to #t if the overflow toggle is set, and to #f otherwise. The value of the overflow toggle can be modified using mix-set-over!.

[Function] mix-cmp

[Function] mix-set-cmp! ’L|’E|’G

Evaluate and set the comparison flag. Possible values are the scheme symbols L (lesser), E (equal) and G (greater).

[Function] mix-up-time

Evaluates to the current virtual machine uptime.

[Function] mix-lap-time

Evaluates to the current virtual machine lapsed time, i.e., the time elapsed since the last run or next command.

[Function] mix-prog-time

Evaluates to the total time spent executing the currently loaded program.

[Function] mix-prog-name

Evaluates to a string containing the basename (without any leading path) of the currently loaded MIX program.

[Function] mix-prog-path

Evaluates to a string containing the full path to the currently loaded MIX program.

[Function] mix-src-path

Evaluates to a string containing the full path to the source file of the currently loaded MIX program.

[Function] mix-src-line [lineno]

[Function] mix-src-line-no

mix-src-line-no evaluates to the current source file number during the execution of a program. mix-src-line evaluates to a string containing the source file line number lineno; when invoked without argument, it evaluates to (mix-src-line (mix-src-

line-no)).

[Function] mix-ddir

Evaluates to a string containing the full path of the current device directory.

9 Reporting Bugs

If you have any questions, comments or suggestions, please send electronic mail to the

author (jao@gnu.org).

If you find a bug in mdk, please send electronic mail to the mdk bug list (bug-mdk@gnu.

org).

In your report, please include the version number, which you can find by running ‘mixasm --version’. Also include in your message the output that the program produced and the output you expected.

Appendix A Copying

GNU MDK is distributed under the GNU General Public License (GPL) and this manual under the GNU Free Documentation License (GFDL).

A.1 GNU General Public License

Version 3, 29 June 2007

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2007 Free Software Foundation, Inc. http://fsf.org/

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

Preamble

The GNU General Public License is a free, copyleft license for software and other kinds of works.

The licenses for most software and other practical works are designed to take away your freedom to share and change the works. By contrast, the GNU General Public License is intended to guarantee your freedom to share and change all versions of a program—to make sure it remains free software for all its users. We, the Free Software Foundation, use the GNU General Public License for most of our software; it applies also to any other work released this way by its authors. You can apply it to your programs, too.

When we speak of free software, we are referring to freedom, not price. Our General Public Licenses are designed to make sure that you have the freedom to distribute copies of free software (and charge for them if you wish), that you receive source code or can get it if you want it, that you can change the software or use pieces of it in new free programs, and that you know you can do these things.

To protect your rights, we need to prevent others from denying you these rights or asking you to surrender the rights. Therefore, you have certain responsibilities if you distribute copies of the software, or if you modify it: responsibilities to respect the freedom of others.

For example, if you distribute copies of such a program, whether gratis or for a fee, you must pass on to the recipients the same freedoms that you received. You must make sure that they, too, receive or can get the source code. And you must show them these terms so they know their rights.

Developers that use the GNU GPL protect your rights with two steps: (1) assert copyright on the software, and (2) offer you this License giving you legal permission to copy, distribute and/or modify it.

For the developers’ and authors’ protection, the GPL clearly explains that there is no warranty for this free software. For both users’ and authors’ sake, the GPL requires that modified versions be marked as changed, so that their problems will not be attributed erroneously to authors of previous versions.

Some devices are designed to deny users access to install or run modified versions of the software inside them, although the manufacturer can do so. This is fundamentally incompatible with the aim of protecting users’ freedom to change the software. The systematic pattern of such abuse occurs in the area of products for individuals to use, which is precisely where it is most unacceptable. Therefore, we have designed this version of the GPL to prohibit the practice for those products. If such problems arise substantially in other domains, we stand ready to extend this provision to those domains in future versions of the GPL, as needed to protect the freedom of users.

Finally, every program is threatened constantly by software patents. States should not allow patents to restrict development and use of software on general-purpose computers, but in those that do, we wish to avoid the special danger that patents applied to a free program could make it effectively proprietary. To prevent this, the GPL assures that patents cannot be used to render the program non-free.

The precise terms and conditions for copying, distribution and modification follow.

TERMS AND CONDITIONS

0. Definitions. “This License” refers to version 3 of the GNU General Public License. “Copyright” also means copyright-like laws that apply to other kinds of works, such as semiconductor masks. “The Program” refers to any copyrightable work licensed under this License. Each licensee is addressed as “you”. “Licensees” and “recipients” may be individuals or organizations. To “modify” a work means to copy from or adapt all or part of the work in a fashion requiring copyright permission, other than the making of an exact copy. The resulting work is called a “modified version” of the earlier work or a work “based on” the earlier work. A “covered work” means either the unmodified Program or a work based on the Program. To “propagate” a work means to do anything with it that, without permission, would make you directly or secondarily liable for infringement under applicable copyright law, except executing it on a computer or modifying a private copy. Propagation includes copying, distribution (with or without modification), making available to the public, and in some countries other activities as well. To “convey” a work means any kind of propagation that enables other parties to make or receive copies. Mere interaction with a user through a computer network, with no transfer of a copy, is not conveying. An interactive user interface displays “Appropriate Legal Notices” to the extent that it includes a convenient and prominently visible feature that (1) displays an appropriate copyright notice, and (2) tells the user that there is no warranty for the work (except to the extent that warranties are provided), that licensees may convey the work under this License, and how to view a copy of this License. If the interface presents a list of user commands or options, such as a menu, a prominent item in the list meets this criterion.

1. Source Code. The “source code” for a work means the preferred form of the work for making modifications to it. “Object code” means any non-source form of a work.

A “Standard Interface” means an interface that either is an official standard defined by a recognized standards body, or, in the case of interfaces specified for a particular programming language, one that is widely used among developers working in that language.

The “System Libraries” of an executable work include anything, other than the work as a whole, that (a) is included in the normal form of packaging a Major Component, but which is not part of that Major Component, and (b) serves only to enable use of the work with that Major Component, or to implement a Standard Interface for which an implementation is available to the public in source code form. A “Major Component”, in this context, means a major essential component (kernel, window system, and so on) of the specific operating system (if any) on which the executable work runs, or a compiler used to produce the work, or an object code interpreter used to run it.

The “Corresponding Source” for a work in object code form means all the source code needed to generate, install, and (for an executable work) run the object code and to modify the work, including scripts to control those activities. However, it does not include the work’s System Libraries, or general-purpose tools or generally available free programs which are used unmodified in performing those activities but which are not part of the work. For example, Corresponding Source includes interface definition files associated with source files for the work, and the source code for shared libraries and dynamically linked subprograms that the work is specifically designed to require, such as by intimate data communication or control flow between those subprograms and other parts of the work.

The Corresponding Source need not include anything that users can regenerate automatically from other parts of the Corresponding Source.

The Corresponding Source for a work in source code form is that same work.

2. Basic Permissions.

All rights granted under this License are granted for the term of copyright on the Program, and are irrevocable provided the stated conditions are met. This License explicitly affirms your unlimited permission to run the unmodified Program. The output from running a covered work is covered by this License only if the output, given its content, constitutes a covered work. This License acknowledges your rights of fair use or other equivalent, as provided by copyright law.

You may make, run and propagate covered works that you do not convey, without conditions so long as your license otherwise remains in force. You may convey covered works to others for the sole purpose of having them make modifications exclusively for you, or provide you with facilities for running those works, provided that you comply with the terms of this License in conveying all material for which you do not control copyright. Those thus making or running the covered works for you must do so exclusively on your behalf, under your direction and control, on terms that prohibit them from making any copies of your copyrighted material outside their relationship with you.

Conveying under any other circumstances is permitted solely under the conditions stated below. Sublicensing is not allowed; section 10 makes it unnecessary.

3. Protecting Users’ Legal Rights From Anti-Circumvention Law.

No covered work shall be deemed part of an effective technological measure under any applicable law fulfilling obligations under article 11 of the WIPO copyright treaty adopted on 20 December 1996, or similar laws prohibiting or restricting circumvention of such measures. When you convey a covered work, you waive any legal power to forbid circumvention of technological measures to the extent such circumvention is effected by exercising rights under this License with respect to the covered work, and you disclaim any intention to limit operation or modification of the work as a means of enforcing, against the work’s users, your or third parties’ legal rights to forbid circumvention of technological measures.

4. Conveying Verbatim Copies.

You may convey verbatim copies of the Program’s source code as you receive it, in any medium, provided that you conspicuously and appropriately publish on each copy an appropriate copyright notice; keep intact all notices stating that this License and any non-permissive terms added in accord with section 7 apply to the code; keep intact all notices of the absence of any warranty; and give all recipients a copy of this License along with the Program. You may charge any price or no price for each copy that you convey, and you may offer support or warranty protection for a fee.

5. Conveying Modified Source Versions. You may convey a work based on the Program, or the modifications to produce it from the Program, in the form of source code under the terms of section 4, provided that you also meet all of these conditions: a. The work must carry prominent notices stating that you modified it, and giving a

relevant date. b. The work must carry prominent notices stating that it is released under this Li-

cense and any conditions added under section 7. This requirement modifies the requirement in section 4 to “keep intact all notices”.

c. You must license the entire work, as a whole, under this License to anyone who comes into possession of a copy. This License will therefore apply, along with any applicable section 7 additional terms, to the whole of the work, and all its parts, regardless of how they are packaged. This License gives no permission to license the work in any other way, but it does not invalidate such permission if you have separately received it.

d. If the work has interactive user interfaces, each must display Appropriate Legal Notices; however, if the Program has interactive interfaces that do not display Appropriate Legal Notices, your work need not make them do so.

A compilation of a covered work with other separate and independent works, which are not by their nature extensions of the covered work, and which are not combined with it such as to form a larger program, in or on a volume of a storage or distribution medium, is called an “aggregate” if the compilation and its resulting copyright are not used to limit the access or legal rights of the compilation’s users beyond what the individual works permit. Inclusion of a covered work in an aggregate does not cause this License to apply to the other parts of the aggregate.

6. Conveying Non-Source Forms.

You may convey a covered work in object code form under the terms of sections 4 and 5, provided that you also convey the machine-readable Corresponding Source under the terms of this License, in one of these ways:

a. Convey the object code in, or embodied in, a physical product (including a physical distribution medium), accompanied by the Corresponding Source fixed on a durable physical medium customarily used for software interchange.

b. Convey the object code in, or embodied in, a physical product (including a physical distribution medium), accompanied by a written offer, valid for at least three years and valid for as long as you offer spare parts or customer support for that product model, to give anyone who possesses the object code either (1) a copy of the Corresponding Source for all the software in the product that is covered by this License, on a durable physical medium customarily used for software interchange, for a price no more than your reasonable cost of physically performing this conveying of source, or (2) access to copy the Corresponding Source from a network server at no charge.

c. Convey individual copies of the object code with a copy of the written offer to provide the Corresponding Source. This alternative is allowed only occasionally and noncommercially, and only if you received the object code with such an offer, in accord with subsection 6b.

d. Convey the object code by offering access from a designated place (gratis or for a charge), and offer equivalent access to the Corresponding Source in the same way through the same place at no further charge. You need not require recipients to copy the Corresponding Source along with the object code. If the place to copy the object code is a network server, the Corresponding Source may be on a different server (operated by you or a third party) that supports equivalent copying facilities, provided you maintain clear directions next to the object code saying where to find the Corresponding Source. Regardless of what server hosts the Corresponding Source, you remain obligated to ensure that it is available for as long as needed to satisfy these requirements.

e. Convey the object code using peer-to-peer transmission, provided you inform other peers where the object code and Corresponding Source of the work are being offered to the general public at no charge under subsection 6d.

A separable portion of the object code, whose source code is excluded from the Corresponding Source as a System Library, need not be included in conveying the object code work.

A “User Product” is either (1) a “consumer product”, which means any tangible personal property which is normally used for personal, family, or household purposes, or (2) anything designed or sold for incorporation into a dwelling. In determining whether a product is a consumer product, doubtful cases shall be resolved in favor of coverage. For a particular product received by a particular user, “normally used” refers to a typical or common use of that class of product, regardless of the status of the particular user or of the way in which the particular user actually uses, or expects or is expected to use, the product. A product is a consumer product regardless of whether the product has substantial commercial, industrial or non-consumer uses, unless such uses represent the only significant mode of use of the product.

“Installation Information” for a User Product means any methods, procedures, authorization keys, or other information required to install and execute modified versions of a covered work in that User Product from a modified version of its Corresponding Source. The information must suffice to ensure that the continued functioning of the modified object code is in no case prevented or interfered with solely because modification has been made.

If you convey an object code work under this section in, or with, or specifically for use in, a User Product, and the conveying occurs as part of a transaction in which the right of possession and use of the User Product is transferred to the recipient in perpetuity or for a fixed term (regardless of how the transaction is characterized), the Corresponding Source conveyed under this section must be accompanied by the Installation Information. But this requirement does not apply if neither you nor any third party retains the ability to install modified object code on the User Product (for example, the work has been installed in ROM).

The requirement to provide Installation Information does not include a requirement to continue to provide support service, warranty, or updates for a work that has been modified or installed by the recipient, or for the User Product in which it has been modified or installed. Access to a network may be denied when the modification itself materially and adversely affects the operation of the network or violates the rules and protocols for communication across the network.

Corresponding Source conveyed, and Installation Information provided, in accord with this section must be in a format that is publicly documented (and with an implementation available to the public in source code form), and must require no special password or key for unpacking, reading or copying.

7. Additional Terms.

“Additional permissions” are terms that supplement the terms of this License by making exceptions from one or more of its conditions. Additional permissions that are applicable to the entire Program shall be treated as though they were included in this License, to the extent that they are valid under applicable law. If additional permissions apply only to part of the Program, that part may be used separately under those permissions, but the entire Program remains governed by this License without regard to the additional permissions.

When you convey a copy of a covered work, you may at your option remove any additional permissions from that copy, or from any part of it. (Additional permissions may be written to require their own removal in certain cases when you modify the work.) You may place additional permissions on material, added by you to a covered work, for which you have or can give appropriate copyright permission.

Notwithstanding any other provision of this License, for material you add to a covered work, you may (if authorized by the copyright holders of that material) supplement the terms of this License with terms:

a. Disclaiming warranty or limiting liability differently from the terms of sections 15 and 16 of this License; or b. Requiring preservation of specified reasonable legal notices or author attributions

in that material or in the Appropriate Legal Notices displayed by works containing it; or

c. Prohibiting misrepresentation of the origin of that material, or requiring that modified versions of such material be marked in reasonable ways as different from the original version; or

d. Limiting the use for publicity purposes of names of licensors or authors of the

material; or e. Declining to grant rights under trademark law for use of some trade names, trade-

marks, or service marks; or f. Requiring indemnification of licensors and authors of that material by anyone who

conveys the material (or modified versions of it) with contractual assumptions of liability to the recipient, for any liability that these contractual assumptions directly impose on those licensors and authors.

All other non-permissive additional terms are considered “further restrictions” within the meaning of section 10. If the Program as you received it, or any part of it, contains a notice stating that it is governed by this License along with a term that is a further restriction, you may remove that term. If a license document contains a further restriction but permits relicensing or conveying under this License, you may add to a covered work material governed by the terms of that license document, provided that the further restriction does not survive such relicensing or conveying. If you add terms to a covered work in accord with this section, you must place, in the relevant source files, a statement of the additional terms that apply to those files, or a notice indicating where to find the applicable terms. Additional terms, permissive or non-permissive, may be stated in the form of a separately written license, or stated as exceptions; the above requirements apply either way.

8. Termination. You may not propagate or modify a covered work except as expressly provided under this License. Any attempt otherwise to propagate or modify it is void, and will automatically terminate your rights under this License (including any patent licenses granted under the third paragraph of section 11). However, if you cease all violation of this License, then your license from a particular copyright holder is reinstated (a) provisionally, unless and until the copyright holder explicitly and finally terminates your license, and (b) permanently, if the copyright holder fails to notify you of the violation by some reasonable means prior to 60 days after the cessation. Moreover, your license from a particular copyright holder is reinstated permanently if the copyright holder notifies you of the violation by some reasonable means, this is the first time you have received notice of violation of this License (for any work) from that copyright holder, and you cure the violation prior to 30 days after your receipt of the notice. Termination of your rights under this section does not terminate the licenses of parties who have received copies or rights from you under this License. If your rights have been terminated and not permanently reinstated, you do not qualify to receive new licenses for the same material under section 10.

9. Acceptance Not Required for Having Copies. You are not required to accept this License in order to receive or run a copy of the Program. Ancillary propagation of a covered work occurring solely as a consequence of using peer-to-peer transmission to receive a copy likewise does not require acceptance. However, nothing other than this License grants you permission to propagate or modify any covered work. These actions infringe copyright if you do not accept this License. Therefore, by modifying or propagating a covered work, you indicate your acceptance of this License to do so.

10. Automatic Licensing of Downstream Recipients. Each time you convey a covered work, the recipient automatically receives a license from the original licensors, to run, modify and propagate that work, subject to this License. You are not responsible for enforcing compliance by third parties with this License. An “entity transaction” is a transaction transferring control of an organization, or substantially all assets of one, or subdividing an organization, or merging organizations. If propagation of a covered work results from an entity transaction, each party to that transaction who receives a copy of the work also receives whatever licenses to the work the party’s predecessor in interest had or could give under the previous paragraph, plus a right to possession of the Corresponding Source of the work from the predecessor in interest, if the predecessor has it or can get it with reasonable efforts. You may not impose any further restrictions on the exercise of the rights granted or affirmed under this License. For example, you may not impose a license fee, royalty, or other charge for exercise of rights granted under this License, and you may not initiate litigation (including a cross-claim or counterclaim in a lawsuit) alleging that any patent claim is infringed by making, using, selling, offering for sale, or importing the Program or any portion of it.

11. Patents.

A “contributor” is a copyright holder who authorizes use under this License of the Program or a work on which the Program is based. The work thus licensed is called the contributor’s “contributor version”. A contributor’s “essential patent claims” are all patent claims owned or controlled by the contributor, whether already acquired or hereafter acquired, that would be infringed by some manner, permitted by this License, of making, using, or selling its contributor version, but do not include claims that would be infringed only as a consequence of further modification of the contributor version. For purposes of this definition, “control” includes the right to grant patent sublicenses in a manner consistent with the requirements of this License. Each contributor grants you a non-exclusive, worldwide, royalty-free patent license under the contributor’s essential patent claims, to make, use, sell, offer for sale, import and otherwise run, modify and propagate the contents of its contributor version. In the following three paragraphs, a “patent license” is any express agreement or commitment, however denominated, not to enforce a patent (such as an express permission to practice a patent or covenant not to sue for patent infringement). To “grant” such a patent license to a party means to make such an agreement or commitment not to enforce a patent against the party. If you convey a covered work, knowingly relying on a patent license, and the Corresponding Source of the work is not available for anyone to copy, free of charge and under the terms of this License, through a publicly available network server or other readily accessible means, then you must either (1) cause the Corresponding Source to be so available, or (2) arrange to deprive yourself of the benefit of the patent license for this particular work, or (3) arrange, in a manner consistent with the requirements of this License, to extend the patent license to downstream recipients. “Knowingly relying” means you have actual knowledge that, but for the patent license, your conveying the covered work in a country, or your recipient’s use of the covered work in a country, would infringe one or more identifiable patents in that country that you have reason to believe are valid. If, pursuant to or in connection with a single transaction or arrangement, you convey, or propagate by procuring conveyance of, a covered work, and grant a patent license to some of the parties receiving the covered work authorizing them to use, propagate, modify or convey a specific copy of the covered work, then the patent license you grant is automatically extended to all recipients of the covered work and works based on it. A patent license is “discriminatory” if it does not include within the scope of its coverage, prohibits the exercise of, or is conditioned on the non-exercise of one or more of the rights that are specifically granted under this License. You may not convey a covered work if you are a party to an arrangement with a third party that is in the business of distributing software, under which you make payment to the third party based on the extent of your activity of conveying the work, and under which the third party grants, to any of the parties who would receive the covered work from you, a discriminatory patent license (a) in connection with copies of the covered work conveyed by you (or copies made from those copies), or (b) primarily for and in connection with specific products or compilations that contain the covered work, unless you entered into that arrangement, or that patent license was granted, prior to 28 March 2007. Nothing in this License shall be construed as excluding or limiting any implied license or other defenses to infringement that may otherwise be available to you under applicable patent law.

12. No Surrender of Others’ Freedom.

If conditions are imposed on you (whether by court order, agreement or otherwise) that contradict the conditions of this License, they do not excuse you from the conditions of this License. If you cannot convey a covered work so as to satisfy simultaneously your obligations under this License and any other pertinent obligations, then as a consequence you may not convey it at all. For example, if you agree to terms that obligate you to collect a royalty for further conveying from those to whom you convey the Program, the only way you could satisfy both those terms and this License would be to refrain entirely from conveying the Program.

13. Use with the GNU Affero General Public License. Notwithstanding any other provision of this License, you have permission to link or combine any covered work with a work licensed under version 3 of the GNU Affero General Public License into a single combined work, and to convey the resulting work.

The terms of this License will continue to apply to the part which is the covered work, but the special requirements of the GNU Affero General Public License, section 13, concerning interaction through a network will apply to the combination as such.

14. Revised Versions of this License. The Free Software Foundation may publish revised and/or new versions of the GNU General Public License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns. Each version is given a distinguishing version number. If the Program specifies that a certain numbered version of the GNU General Public License “or any later version” applies to it, you have the option of following the terms and conditions either of that numbered version or of any later version published by the Free Software Foundation. If the Program does not specify a version number of the GNU General Public License, you may choose any version ever published by the Free Software Foundation. If the Program specifies that a proxy can decide which future versions of the GNU General Public License can be used, that proxy’s public statement of acceptance of a version permanently authorizes you to choose that version for the Program. Later license versions may give you additional or different permissions. However, no additional obligations are imposed on any author or copyright holder as a result of your choosing to follow a later version.

15. Disclaimer of Warranty.

THERE IS NO WARRANTY FOR THE PROGRAM, TO THE EXTENT PER-MITTED BY APPLICABLE LAW. EXCEPT WHEN OTHERWISE STATED IN WRITING THE COPYRIGHT HOLDERS AND/OR OTHER PARTIES PROVIDE THE PROGRAM “AS IS” WITHOUT WARRANTY OF ANY KIND, EITHER EX-PRESSED OR IMPLIED, INCLUDING, BUT NOT LIMITED TO, THE IMPLIED WARRANTIES OF MERCHANTABILITY AND FITNESS FOR A PARTICULAR PURPOSE. THE ENTIRE RISK AS TO THE QUALITY AND PERFORMANCE OF THE PROGRAM IS WITH YOU. SHOULD THE PROGRAM PROVE DEFEC-TIVE, YOU ASSUME THE COST OF ALL NECESSARY SERVICING, REPAIR OR CORRECTION.

16. Limitation of Liability.

IN NO EVENT UNLESS REQUIRED BY APPLICABLE LAW OR AGREED TO IN WRITING WILL ANY COPYRIGHT HOLDER, OR ANY OTHER PARTY WHO MODIFIES AND/OR CONVEYS THE PROGRAM AS PERMITTED ABOVE, BE LIABLE TO YOU FOR DAMAGES, INCLUDING ANY GENERAL, SPECIAL, IN-CIDENTAL OR CONSEQUENTIAL DAMAGES ARISING OUT OF THE USE OR INABILITY TO USE THE PROGRAM (INCLUDING BUT NOT LIMITED TO LOSS OF DATA OR DATA BEING RENDERED INACCURATE OR LOSSES SUS-TAINED BY YOU OR THIRD PARTIES OR A FAILURE OF THE PROGRAM TO OPERATE WITH ANY OTHER PROGRAMS), EVEN IF SUCH HOLDER OR OTHER PARTY HAS BEEN ADVISED OF THE POSSIBILITY OF SUCH DAM-AGES.

17. Interpretation of Sections 15 and 16. If the disclaimer of warranty and limitation of liability provided above cannot be given local legal effect according to their terms, reviewing courts shall apply local law that most closely approximates an absolute waiver of all civil liability in connection with the Program, unless a warranty or assumption of liability accompanies a copy of the Program in return for a fee.

END OF TERMS AND CONDITIONS

How to Apply These Terms to Your New Programs

If you develop a new program, and you want it to be of the greatest possible use to the public, the best way to achieve this is to make it free software which everyone can redistribute and change under these terms.

To do so, attach the following notices to the program. It is safest to attach them to the start of each source file to most effectively state the exclusion of warranty; and each file should have at least the “copyright” line and a pointer to where the full notice is found.

      one line to give the program's name and a brief idea of what it does.
      Copyright (C) year name of author
      This program is free software: you can redistribute it and/or modify
      it under the terms of the GNU General Public License as published by
      the Free Software Foundation, either version 3 of the License, or (at
      your option) any later version.
      This program is distributed in the hope that it will be useful, but
      WITHOUT ANY WARRANTY; without even the implied warranty of
      MERCHANTABILITY or FITNESS FOR A PARTICULAR PURPOSE.  See the GNU
      General Public License for more details.
      You should have received a copy of the GNU General Public License
      along with this program.  If not, see http://www.gnu.org/licenses/.

Also add information on how to contact you by electronic and paper mail.

If the program does terminal interaction, make it output a short notice like this when it starts in an interactive mode:

      program Copyright (C) year name of author
      This program comes with ABSOLUTELY NO WARRANTY; for details type `show w'.
      This is free software, and you are welcome to redistribute it
      under certain conditions; type `show c' for details.

The hypothetical commands ‘show w’ and ‘show c’ should show the appropriate parts of the General Public License. Of course, your program’s commands might be different; for a GUI interface, you would use an “about box”.

You should also get your employer (if you work as a programmer) or school, if any, to sign a “copyright disclaimer” for the program, if necessary. For more information on this, and how to apply and follow the GNU GPL, see http://www.gnu.org/licenses/.

The GNU General Public License does not permit incorporating your program into proprietary programs. If your program is a subroutine library, you may consider it more useful to permit linking proprietary applications with the library. If this is what you want to do, use the GNU Lesser General Public License instead of this License. But first, please

read http://www.gnu.org/philosophy/why-not-lgpl.html.

A.2 GNU Free Documentation License

Version 1.2, November 2002

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2000,2001,2002 Free Software Foundation, Inc. 51 Franklin St, Fifth Floor, Boston, MA 02110-1301, USA

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

0. PREAMBLE The purpose of this License is to make a manual, textbook, or other functional and useful document free in the sense of freedom: to assure everyone the effective freedom to copy and redistribute it, with or without modifying it, either commercially or noncommercially. Secondarily, this License preserves for the author and publisher a way to get credit for their work, while not being considered responsible for modifications made by others. This License is a kind of “copyleft”, which means that derivative works of the document must themselves be free in the same sense. It complements the GNU General Public License, which is a copyleft license designed for free software. We have designed this License in order to use it for manuals for free software, because free software needs free documentation: a free program should come with manuals providing the same freedoms that the software does. But this License is not limited to software manuals; it can be used for any textual work, regardless of subject matter or whether it is published as a printed book. We recommend this License principally for works whose purpose is instruction or reference.

1. APPLICABILITY AND DEFINITIONS This License applies to any manual or other work, in any medium, that contains a notice placed by the copyright holder saying it can be distributed under the terms of this License. Such a notice grants a world-wide, royalty-free license, unlimited in duration, to use that work under the conditions stated herein. The “Document”, below, refers to any such manual or work. Any member of the public is a licensee, and is addressed as “you”. You accept the license if you copy, modify or distribute the work in a way requiring permission under copyright law. A “Modified Version” of the Document means any work containing the Document or a portion of it, either copied verbatim, or with modifications and/or translated into another language. A “Secondary Section” is a named appendix or a front-matter section of the Document that deals exclusively with the relationship of the publishers or authors of the Document to the Document’s overall subject (or to related matters) and contains nothing that could fall directly within that overall subject. (Thus, if the Document is in part a textbook of mathematics, a Secondary Section may not explain any mathematics.) The relationship could be a matter of historical connection with the subject or with related matters, or of legal, commercial, philosophical, ethical or political position regarding them. The “Invariant Sections” are certain Secondary Sections whose titles are designated, as being those of Invariant Sections, in the notice that says that the Document is released under this License. If a section does not fit the above definition of Secondary then it is not allowed to be designated as Invariant. The Document may contain zero Invariant Sections. If the Document does not identify any Invariant Sections then there are none. The “Cover Texts” are certain short passages of text that are listed, as Front-Cover Texts or Back-Cover Texts, in the notice that says that the Document is released under this License. A Front-Cover Text may be at most 5 words, and a Back-Cover Text may be at most 25 words. A “Transparent” copy of the Document means a machine-readable copy, represented in a format whose specification is available to the general public, that is suitable for revising the document straightforwardly with generic text editors or (for images composed of pixels) generic paint programs or (for drawings) some widely available drawing editor, and that is suitable for input to text formatters or for automatic translation to a variety of formats suitable for input to text formatters. A copy made in an otherwise Transparent file format whose markup, or absence of markup, has been arranged to thwart or discourage subsequent modification by readers is not Transparent. An image format is not Transparent if used for any substantial amount of text. A copy that is not “Transparent” is called “Opaque”. Examples of suitable formats for Transparent copies include plain ascii without markup, Texinfo input format, LaTEX input format, SGML or XML using a publicly available DTD, and standard-conforming simple HTML, PostScript or PDF designed for human modification. Examples of transparent image formats include PNG, XCF and JPG. Opaque formats include proprietary formats that can be read and edited only by proprietary word processors, SGML or XML for which the DTD and/or processing tools are not generally available, and the machine-generated HTML, PostScript or PDF produced by some word processors for output purposes only. The “Title Page” means, for a printed book, the title page itself, plus such following pages as are needed to hold, legibly, the material this License requires to appear in the title page. For works in formats which do not have any title page as such, “Title Page” means the text near the most prominent appearance of the work’s title, preceding the beginning of the body of the text. A section “Entitled XYZ” means a named subunit of the Document whose title either is precisely XYZ or contains XYZ in parentheses following text that translates XYZ in another language. (Here XYZ stands for a specific section name mentioned below, such as “Acknowledgements”, “Dedications”, “Endorsements”, or “History”.) To “Preserve the Title” of such a section when you modify the Document means that it remains a section “Entitled XYZ” according to this definition. The Document may include Warranty Disclaimers next to the notice which states that this License applies to the Document. These Warranty Disclaimers are considered to be included by reference in this License, but only as regards disclaiming warranties: any other implication that these Warranty Disclaimers may have is void and has no effect on the meaning of this License.

2. VERBATIM COPYING You may copy and distribute the Document in any medium, either commercially or noncommercially, provided that this License, the copyright notices, and the license notice saying this License applies to the Document are reproduced in all copies, and that you add no other conditions whatsoever to those of this License. You may not use technical measures to obstruct or control the reading or further copying of the copies you make or distribute. However, you may accept compensation in exchange for copies. If you distribute a large enough number of copies you must also follow the conditions in section 3.

You may also lend copies, under the same conditions stated above, and you may publicly display copies.

3. COPYING IN QUANTITY

If you publish printed copies (or copies in media that commonly have printed covers) of the Document, numbering more than 100, and the Document’s license notice requires Cover Texts, you must enclose the copies in covers that carry, clearly and legibly, all these Cover Texts: Front-Cover Texts on the front cover, and Back-Cover Texts on the back cover. Both covers must also clearly and legibly identify you as the publisher of these copies. The front cover must present the full title with all words of the title equally prominent and visible. You may add other material on the covers in addition. Copying with changes limited to the covers, as long as they preserve the title of the Document and satisfy these conditions, can be treated as verbatim copying in other respects.

If the required texts for either cover are too voluminous to fit legibly, you should put the first ones listed (as many as fit reasonably) on the actual cover, and continue the rest onto adjacent pages.

If you publish or distribute Opaque copies of the Document numbering more than 100, you must either include a machine-readable Transparent copy along with each Opaque copy, or state in or with each Opaque copy a computer-network location from which the general network-using public has access to download using public-standard network protocols a complete Transparent copy of the Document, free of added material. If you use the latter option, you must take reasonably prudent steps, when you begin distribution of Opaque copies in quantity, to ensure that this Transparent copy will remain thus accessible at the stated location until at least one year after the last time you distribute an Opaque copy (directly or through your agents or retailers) of that edition to the public.

It is requested, but not required, that you contact the authors of the Document well before redistributing any large number of copies, to give them a chance to provide you with an updated version of the Document.

4. MODIFICATIONS

You may copy and distribute a Modified Version of the Document under the conditions of sections 2 and 3 above, provided that you release the Modified Version under precisely this License, with the Modified Version filling the role of the Document, thus licensing distribution and modification of the Modified Version to whoever possesses a copy of it. In addition, you must do these things in the Modified Version:

A. Use in the Title Page (and on the covers, if any) a title distinct from that of the Document, and from those of previous versions (which should, if there were any, be listed in the History section of the Document). You may use the same title as a previous version if the original publisher of that version gives permission.

B. List on the Title Page, as authors, one or more persons or entities responsible for authorship of the modifications in the Modified Version, together with at least five of the principal authors of the Document (all of its principal authors, if it has fewer than five), unless they release you from this requirement.

C. State on the Title page the name of the publisher of the Modified Version, as the

publisher. D. Preserve all the copyright notices of the Document. E. Add an appropriate copyright notice for your modifications adjacent to the other

copyright notices. F. Include, immediately after the copyright notices, a license notice giving the public

permission to use the Modified Version under the terms of this License, in the form shown in the Addendum below.

G. Preserve in that license notice the full lists of Invariant Sections and required Cover

Texts given in the Document’s license notice. H. Include an unaltered copy of this License. I. Preserve the section Entitled “History”, Preserve its Title, and add to it an item

stating at least the title, year, new authors, and publisher of the Modified Version as given on the Title Page. If there is no section Entitled “History” in the Document, create one stating the title, year, authors, and publisher of the Document as given on its Title Page, then add an item describing the Modified Version as stated in the previous sentence.

J. Preserve the network location, if any, given in the Document for public access to a Transparent copy of the Document, and likewise the network locations given in the Document for previous versions it was based on. These may be placed in the “History” section. You may omit a network location for a work that was published at least four years before the Document itself, or if the original publisher of the version it refers to gives permission.

K. For any section Entitled “Acknowledgements” or “Dedications”, Preserve the Title of the section, and preserve in the section all the substance and tone of each of the contributor acknowledgements and/or dedications given therein.

L. Preserve all the Invariant Sections of the Document, unaltered in their text and in their titles. Section numbers or the equivalent are not considered part of the section titles.

M. Delete any section Entitled “Endorsements”. Such a section may not be included

in the Modified Version. N. Do not retitle any existing section to be Entitled “Endorsements” or to conflict in

title with any Invariant Section. O. Preserve any Warranty Disclaimers.

If the Modified Version includes new front-matter sections or appendices that qualify as Secondary Sections and contain no material copied from the Document, you may at your option designate some or all of these sections as invariant. To do this, add their titles to the list of Invariant Sections in the Modified Version’s license notice. These titles must be distinct from any other section titles.

You may add a section Entitled “Endorsements”, provided it contains nothing but endorsements of your Modified Version by various parties—for example, statements of peer review or that the text has been approved by an organization as the authoritative definition of a standard. You may add a passage of up to five words as a Front-Cover Text, and a passage of up to 25 words as a Back-Cover Text, to the end of the list of Cover Texts in the Modified Version. Only one passage of Front-Cover Text and one of Back-Cover Text may be added by (or through arrangements made by) any one entity. If the Document already includes a cover text for the same cover, previously added by you or by arrangement made by the same entity you are acting on behalf of, you may not add another; but you may replace the old one, on explicit permission from the previous publisher that added the old one. The author(s) and publisher(s) of the Document do not by this License give permission to use their names for publicity for or to assert or imply endorsement of any Modified Version.

5. COMBINING DOCUMENTS You may combine the Document with other documents released under this License, under the terms defined in section 4 above for modified versions, provided that you include in the combination all of the Invariant Sections of all of the original documents, unmodified, and list them all as Invariant Sections of your combined work in its license notice, and that you preserve all their Warranty Disclaimers. The combined work need only contain one copy of this License, and multiple identical Invariant Sections may be replaced with a single copy. If there are multiple Invariant Sections with the same name but different contents, make the title of each such section unique by adding at the end of it, in parentheses, the name of the original author or publisher of that section if known, or else a unique number. Make the same adjustment to the section titles in the list of Invariant Sections in the license notice of the combined work. In the combination, you must combine any sections Entitled “History” in the various original documents, forming one section Entitled “History”; likewise combine any sections Entitled “Acknowledgements”, and any sections Entitled “Dedications”. You must delete all sections Entitled “Endorsements.”

6. COLLECTIONS OF DOCUMENTS You may make a collection consisting of the Document and other documents released under this License, and replace the individual copies of this License in the various documents with a single copy that is included in the collection, provided that you follow the rules of this License for verbatim copying of each of the documents in all other respects. You may extract a single document from such a collection, and distribute it individually under this License, provided you insert a copy of this License into the extracted document, and follow this License in all other respects regarding verbatim copying of that document.

7. AGGREGATION WITH INDEPENDENT WORKS A compilation of the Document or its derivatives with other separate and independent documents or works, in or on a volume of a storage or distribution medium, is called an “aggregate” if the copyright resulting from the compilation is not used to limit the legal rights of the compilation’s users beyond what the individual works permit. When the Document is included in an aggregate, this License does not apply to the other works in the aggregate which are not themselves derivative works of the Document. If the Cover Text requirement of section 3 is applicable to these copies of the Document, then if the Document is less than one half of the entire aggregate, the Document’s Cover Texts may be placed on covers that bracket the Document within the aggregate, or the electronic equivalent of covers if the Document is in electronic form. Otherwise they must appear on printed covers that bracket the whole aggregate.

8. TRANSLATION Translation is considered a kind of modification, so you may distribute translations of the Document under the terms of section 4. Replacing Invariant Sections with translations requires special permission from their copyright holders, but you may include translations of some or all Invariant Sections in addition to the original versions of these Invariant Sections. You may include a translation of this License, and all the license notices in the Document, and any Warranty Disclaimers, provided that you also include the original English version of this License and the original versions of those notices and disclaimers. In case of a disagreement between the translation and the original version of this License or a notice or disclaimer, the original version will prevail. If a section in the Document is Entitled “Acknowledgements”, “Dedications”, or “History”, the requirement (section 4) to Preserve its Title (section 1) will typically require changing the actual title.

9. TERMINATION You may not copy, modify, sublicense, or distribute the Document except as expressly provided for under this License. Any other attempt to copy, modify, sublicense or distribute the Document is void, and will automatically terminate your rights under this License. However, parties who have received copies, or rights, from you under this License will not have their licenses terminated so long as such parties remain in full compliance.

10. FUTURE REVISIONS OF THIS LICENSE The Free Software Foundation may publish new, revised versions of the GNU Free Documentation License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

See http://www.gnu.org/copyleft/.

Each version of the License is given a distinguishing version number. If the Document specifies that a particular numbered version of this License “or any later version” applies to it, you have the option of following the terms and conditions either of that specified version or of any later version that has been published (not as a draft) by the Free Software Foundation. If the Document does not specify a version number of this License, you may choose any version ever published (not as a draft) by the Free Software Foundation.

ADDENDUM: How to use this License for your documents

To use this License in a document you have written, include a copy of the License in the document and put the following copyright and license notices just after the title page:

        Copyright (C)  year  your name.
        Permission is granted to copy, distribute and/or modify this document
        under the terms of the GNU Free Documentation License, Version 1.2
        or any later version published by the Free Software Foundation;
        with no Invariant Sections, no Front-Cover Texts, and no Back-Cover
        Texts.  A copy of the license is included in the section entitled ``GNU
        Free Documentation License''.

If you have Invariant Sections, Front-Cover Texts and Back-Cover Texts, replace the “with…Texts.” line with this:

          with the Invariant Sections being list their titles, with
          the Front-Cover Texts being list, and with the Back-Cover Texts
          being list.

If you have Invariant Sections without Cover Texts, or some other combination of the three, merge those two alternatives to suit the situation.

If your document contains nontrivial examples of program code, we recommend releasing these examples in parallel under your choice of free software license, such as the GNU General Public License, to permit their use in free software.

Concept Index 93

Concept Index

.

.mix file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 .mixal file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 .mls file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

A

address . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 address transfer operators . . . . . . . . . . . . . . . . . . . . . 14 arithmetic operators . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 assembler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28, 45 assembly . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

B

binary operator. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 binary programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 break hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 breakpoints. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 bugs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 byte . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

C

cell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 cm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 comments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 comparison indicator . . . . . . . . . . . . . . . . . . . . . . . . . . 10 comparison operators . . . . . . . . . . . . . . . . . . . . . . . . . . 15 compiling. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 conversion operators . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

E

exection time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

F

FDL, GNU Free Documentation License . . . . . . . 86 field specification . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 fspec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

G

global hook. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 gmixvm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 GPL, GNU General Public License . . . . . . . . . . . . 75 GTK+ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 GUI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59

H

help . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 hook function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36

I

index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 input-output devices . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 input-output operators . . . . . . . . . . . . . . . . . . . . . . . . 16 instruction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 instruction set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 interactive mode . . . . . . . . . . . . . . . . . . . . . . . . . . . 28, 30 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

J

jump operators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

L

literal constants . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 load . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 loading operators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 local symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

M

memory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 memory cell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 miscellaneous operators . . . . . . . . . . . . . . . . . . . . . . . . 18 MIX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 MIX architecture . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 MIX assembly language. . . . . . . . . . . . . . . . . . . . . . . . 19 MIX byte. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 MIX computer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 MIX instruction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 MIX register. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 MIX word . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 MIXAL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9, 19, 27, 45 mixasm. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28, 45 mixguile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33, 65 mixguile options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 mixvm. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28, 47

N

next . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 non-interactive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 non-interactive mode. . . . . . . . . . . . . . . . . . . . . . . 28, 29

O

operator . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 ov . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 overflow toggle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

94 GNU MIX Development Kit (mdk)

P

pc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 pmem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 post-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 pre-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 preg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 problems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 psym . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32

Q

questions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73

R

rA. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 register . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 REPL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 rIn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 rJ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 run. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 rX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

S

sbp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 sbpa . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 Scheme . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 Scheme functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 Scheme script . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 scmf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 shift . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 shift operators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 source file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 storing operators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 suggestions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73

T

tail recursion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 time. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 tutorial. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27

U

un . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 unary operator . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

V

virtual machine. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

W

w-expressions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 word . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

Instructions and commands 95

Instructions and commands

A

About... . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 ADD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 ALF. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

C

cabp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 cbp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 cbpa . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 cbpc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 cbpm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 cbpo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 cbpr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 Change font . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 CHAR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 Clear breakpoints . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 CMPA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 CMPi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 CMPX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 compile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 Compile.... . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 CON. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

D

DECA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 DECi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 DECX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 Detached windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 Device output... . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 Devices dir... . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 DIV. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

E

edit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 Edit... . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 END. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 ENNA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 ENNi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 ENNX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 ENTA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 ENTi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 ENTX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 EQU. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 Exit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 External programs... . . . . . . . . . . . . . . . . . . . . . . . . 63

H

help . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 HLT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

I

IN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 INCA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 INCi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 INCX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 IOC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

J

JAE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 JAN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 JANN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 JANP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 JANZ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 JAO. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 JAP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 JAZ. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 JBUS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 JE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 JG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 JGE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 JiN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 JiNN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 JiNP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 JiNZ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 JiP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 JiZ. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 JL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 JLE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 JMP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 JNE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 JNOV . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 JOV. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 JRED . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 JSJ. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 JXE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 JXN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 JXNN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 JXNP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 JXNZ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 JXO. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 JXP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 JXZ. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

L

LDA. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 LDAN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 LDi. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 LDiN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 LDX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 LDXN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 load . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 Load... . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61

96 GNU MIX Development Kit (mdk)

M

mix-add-break-hook . . . . . . . . . . . . . . . . . . . . . . . . . . 68 mix-add-cond-break . . . . . . . . . . . . . . . . . . . . . . . . . . 68

mix-add-global-post-hook . . . . . . . . . . . . . . . . . . . 68
mix-add-global-pre-hook. . . . . . . . . . . . . . . . . . . . . 68

mix-add-post-hook. . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 mix-add-pre-hook. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 mix-cell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 mix-cmp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 mix-ddir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 mix-lap-time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 mix-last-result . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 mix-loc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 mix-over . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 mix-pmem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 mix-preg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 mix-prog-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 mix-prog-path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 mix-prog-time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 mix-reg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 mix-set-cell! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 mix-set-cmp! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 mix-set-over! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 mix-set-reg! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 mix-sover . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 mix-src-line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 mix-src-line-no . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 mix-src-path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 mix-sreg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 mix-up-time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 mix-vm-break? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 mix-vm-cond-break? . . . . . . . . . . . . . . . . . . . . . . . . . . 69 mix-vm-empty? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 mix-vm-error? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 mix-vm-halted? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 mix-vm-loaded? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 mix-vm-running? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 mix-vm-status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 mixvm-cmd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 mixvm-status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 MOVE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 MUL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

N

next . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 Next . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 NOP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 NUM. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

O

ORIG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 OUT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

P

pall . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 pasm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 pbt. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 pc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 pddir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 pedit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 pflags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 pline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 pmem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 pprog . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 preg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 prompt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 psrc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 pstat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 psym . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 ptime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56

Q

quit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50

R

run. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 Run. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62

S

sasm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 Save . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 Save on exit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 sbp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 sbpc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 sbpm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 sbpo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 sbpr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 scmf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 scmp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 sddir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 sedit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 SLA. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 SLAX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 SLB. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 SLC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 slog . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 smem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 sover . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 spba . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 SRA. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 SRAX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 SRB. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 SRC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 sreg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 ssym . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 STA. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 STi. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 stime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 STJ. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 strace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

Instructions and commands 97

STX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 STZ. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 SUB. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 Symbols.... . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62

T

Toolbar(s). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62

W

w2d. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 weval . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53